### v2

# A mathematical theory of the critical point of ferromagnetic Ising systems

Abstract
We develop a theory of the critical point of the ferromagnetic Ising model, whose basic objects are the ergodic (pure) states of the infinite system. It proves the existence of anomalous critical fluctuations, for dimension $\nu=2$ and, under a standard assumption, for $\nu=3$, for the model with nearest-neighbor interaction, in a way which is consistent with the probabilistic approach of Cassandro, Jona-Lasinio, and several others, reviewed in Jona-Lasinio’s article in Phys. Rep. 352,439 (2001). We propose to single out the state at the critical temperature $T_{c}$, among the ergodic thermal states associated to temperatures $0\leq T\leq T_{c}$, by a condition of non-summable clustering of the connected two-point function. The analogous condition on the connected (2r)- point functions, for $r\geq 2$ , together with a scaling hypothesis, natural within our framework, proves that the (macroscopic) fluctuation state is quasi-free, after a proper rescaling, also at the critical temperature, in agreement with a theorem by Cassandro and Jona-Lasinio, whose proof is, however, shown to be incomplete. Other subjects treated include topics relating to universality, including spontaneous breaking of continuous symmetries and violations of universality in problems of energetic and dynamic stability.
Keywords: critical point; ferromagnetic Ising model; phase transition; criterion of criticality; fluctuation states; universality.
1 General overview on phase transitions, of direct relevance to the present paper
1.1 Introduction and summary
As remarked by Glimm in his review of certain books and articles on statistical mechanics [Gli76], “statistical mechanics concerns infinite systems, and the relevant analysis is over infinite-dimensional spaces”. We shall use his illuminating introdution as a useful guide. Let $\Lambda$ a finite subset of $\mathbb{Z}^{\nu}$, with volume (identified with the number of sites) $|\Lambda|$, with $\nu$ the dimension and $\eta(\Lambda)$ denote the set of nearest-neighbor pairs $x,y\in\Lambda$ with $\left|x-y\right|=1$. Consider the measures on $\mathbb{R}^{\left|\Lambda\right|}$
$d\mu_{T,h}^{\Lambda}\equiv\prod_{x\in\Lambda}\exp(h\sigma_{x}/T)\prod_{x,y\in\eta(\Lambda)}\exp(-J\sigma_{x}\sigma_{y}/T)d\mu_{\Lambda}$ (1.1.1)
where
$d\mu_{\Lambda}\equiv\prod_{x\in\Lambda}d\mu(\sigma_{x})\qquad\text{ with}\qquad\text{ }d\mu(\sigma_{x})=\frac{\delta_{1}(\sigma_{x})+\delta_{-1}(\sigma_{x})}{2}d\sigma_{x}$ (1.1.2)
and $\delta_{a}$ denotes the Dirac delta measure concentrated at $\alpha$ so $\sigma_{x}\in\left\{{-1,1}\right\}$ are discrete Ising variables. $h$ and $J$ denote the real values of the external field and the couplings, and $T$ the absolute temperature (the Boltzmann constant $k$ being set equal to one). $d\mu_{\Lambda}$ defines the model of Ising spins in a finite box $\Lambda$. The free energy $f_{\Lambda}(T,h)$ is defined by
[…]
The thermodynamic limit $\Lambda\nearrow\mathbb{Z}^{\nu}$ (in the simple case of $\Lambda$ a cubic box of side $a$, it is the limit $a\rightarrow\infty$) exists [Thomp72]:
$f(T,h)\equiv\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}f_{\Lambda}(T,h)$ (1.1.4)
While $f_{\Lambda}(T,h)$ is an analytic function of $T$ and $h$ for any $\Lambda$ with $\left|\Lambda\right|<\infty$, $f(T,h)$ is only piecewise analytic, and the boundaries between the analytic pieces of $f(T,h)$ correspond to phase transitions. In the lattice-gas picture with variables $n_{x}=\left(1+\sigma_{x}\right)/2$, with $n_{x}=0\text{ or }1$ according to whether the site $x$ is occupied or empty, it describes, for example, a liquid-vapor transition. In the original variables, there occurs a ferromagnetic phase transition.
As regards the description of real systems, the Ising model in dimension $\nu=3$ is one of the most successful models of both ferromagnets and fluids, due to the fact that most magnetic materials are (highly) anisotropic [PV02], and the lattice-gas analogy is very effective at the critical point [SS09].
The measure $d\mu_{T,h}^{\Lambda}$ is not completely determined by the r.h.s. of equation (1.1.1): we need to impose the boundary conditions (b.c.), where the spins $\sigma_{x}$ at the boundary:
$\ \delta\Lambda\equiv\{x:\text{dist}(x,\Lambda)=1\}$ (1.1.5)
are specified in (1.1.1) with the nearest-neighbors set $\eta(\Lambda)$ replaced by $\eta\left(\Lambda\cup\delta\Lambda\right)$. For high temperature $T$ in (1.1.1),(1.1.2) (“weak couplings” $J/T<<1$), $h=0$, or any value of $J$ if $h\neq 0$),
$d\mu_{T,h}\equiv\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}d\mu_{T,h}^{\Lambda}$ (1.1.6)
is independent of the b.c., but for low temperature $T$ and $h=0$ (“strong coupling”), and $\nu\geq 2$, it is not: this phenomenon is a second notion of a phase transition. A phase transition also occurs in the previously mentioned sense (to which we refer for definiteness as “first notion”) of non-piecewise analyticity. More precisely, $\partial f/\partial h$ has a jump discontinuity (called spontaneous magnetization $m_{s}$) at $h=0$:
$m_{s}\equiv\lim_{h\rightarrow 0+}\frac{\partial f}{\partial h}$ (1.1.7)
Note that the second notion of phase transition is different from the first one, because $f(T,h)$ is independent of the b.c. [Thomp72].
Note also that in (1.1.7), the condition $m_{s}>0$ characterizes a first-order transition, which occurs for $0\leq T<T_{c}$, but, at the critical point, $m_{s}=m_{s}(T)\to 0$ as $T\to T_{c}^{-}$ (also as $T\to T_{c}^{+}$) continuously, see the forthcoming Remark 1.4 and the reference given there. Near $T=T_{c}$, we have the forthcoming (1.1.17), i.e., the susceptibility diverges, and therefore the critical point is a second-order point. This is further confirmed by looking at the manifold of points $(T,h)$, the specific heat diverges at $T=T_{c}$ when $h=0$.
In this paper we restrict ourselves for simplicity to the situation of zero external field ($h=0$): the only free parameter is, therefore, the temperature $T$, which is taken to vary in the range $0\leq T\leq T_{c}$, in which phase transitions occur.
As remarked by Glimm, “Since $d\mu_{T}$ is defined by the limit $\Lambda\nearrow\mathbb{Z}^{\nu}$, $\delta\Lambda\rightarrow\infty$, weak coupling leads to (exact) independence of $\delta\Lambda$ and $d\mu_{T}$, i.e., independence of random variables (r.v.) separated by an infinite distance, while strong coupling allows phase transition and dependence between r.v. separated by an infinite distance”.
In the last half-century, this idea was considerably developed, perhaps most deeply, in the realm of the Ising model, through the probabilistic approach to the theory of the renormalization group (RG), championned by several people, among them M. Cassandro, G. Jona-Lasinio, G. Gallavotti, A. Martin-Löf, P. M. Bleher and Ya. G. Sinai ([CasJL78], [CasJL76], [GML75], [BleSi73], [BleSi75]), and reviewed by Jona-Lasinio in [JL01], to which we refer for numerous additional references. A basic result in this context was Bleher and Sinai’s treatment [BleSi73], [BleSi75] of the critical behavior of Dyson’s hierarchical model [Dy69], which was expected to describe qualitatively well the critical behavior of the Ising model with both short-range and long-range interactions. These seminal papers greatly stimulated further work on hierachical models for classical spin systems, see, in particular, [MCG08] and references therein.
The measure $d\mu_{T}^{\Lambda}=d\mu_{T,h=0}^{\Lambda}$ generates a state $\omega_{T}^{\Lambda}$ over the (abelian) $\mathrm{C}^{\ast}$-algebra of functions $f_{\Lambda}$ of the variables $\{\sigma_{x};\ x\in\Lambda\}$ by
[…]
This is a positive, normalized functional of the $f_{\Lambda}$:
$\omega_{T}^{\Lambda}(f_{\Lambda})\geq 0\qquad\text{ if }\qquad f_{\Lambda}\geq 0\quad\text{ and }\quad\omega_{T}^{\Lambda}(\mathbf{1})=1$ (1.1.9)
The limit measure, the limit $d\mu$ as $\Lambda\nearrow\mathbb{Z}^{\nu}$, $\delta\Lambda\rightarrow\infty$, of the $d\mu_{T}^{\Lambda}$, is associated to a state of the infinite system $\omega_{T}$ by
$\omega_{T}(f)=\!``\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\!"\omega_{T}^{\Lambda}(f_{\Lambda})$ (1.1.10)
where the quotation marks above mean that $f$ is the limit of $f_{\Lambda}$ in the sup norm ([RR67], Section 1). We may also view (this is the natural approach from a physical standpoint) the Ising model as an anisotropic limit of a quantum model: in this case (the forthcoming (1.2.1), one may speak of the “Ising universality class”, for which phase transitions with spontaneous symmetry breaking (s.s.b.) of a discrete symmetry, the spin-flip symmetry $\sigma_{x}\rightarrow-\sigma_{x},~{}\forall x\in\mathbb{Z}^{\nu}$ (and its generalization in the quantum case) take place, see also Subection 1.3.1. The finite-region algebras $\mathcal{A}_{\Lambda}$ are then the spin algebras generated by the Pauli matrices $\{\sigma_{x}^{1,2,3}\}$, and the “quasilocal algebra” $\mathcal{A}$ is the inductive limit of the $\mathcal{A}_{\Lambda}$ (i.e., the norm closure of the union of the $\mathcal{A}_{\Lambda}$, see Subsection 1.2. We adopt this view (which will be particularly useful for the dynamics in Subsection 3.2; restriction to the abelian algebra generated by the $\{\sigma_{x}^{3}\}$ coincides with the previous description.
The study of statistical mechanics from the point of view of the states of infinite systems in both classical and quantum systems has a long history, beginning with the paper of Haag, Hugenholtz and Winnink [HHW67], and the papers of Lanford and Ruelle [LRu69] (see also Ruelle’s book [Rue69]), and going over to the results of uniqueness of the equilibrium state of the Ising model for $T>T_{c}$ due to Dobrushin [D72] and Lebowitz and Martin-Löf [LML72]. The books by Simon [Sim93] and Israel [Isr79] are good sources from the point of view of treating classical and quantum models in parallel.
The dependence vs independence between random variables separated by an infinite distance, which leads to the “sharp dichotomy between strong and weak coupling” mentioned by Glimm ([Gli76], p. 675), “can only exist for infinite systems: in finite systems there is no ’infinitely distant separation’ between the degrees of freedom, i.e., between the coordinate directions in the finite-dimensional space of variables defining the problem”. A natural way to define this dependence precisely is by the rate of clustering of the (necessarily!) states of the infinite system $\omega_{T}$. If the latter are assumed to be translation-invariant, i.e., invariant under the translation $\tau_{x}$ (defined by $\tau_{x}(\sigma_{y}^{3})=\sigma_{x+y}^{3}$), we shall see in Subsection 1.2 that these are “automorphisms of $\mathcal{A}$), this clustering may be expressed by the property (the capital letter “C” below refers to “connected” rather than “critical”): $\forall A,B\in\mathcal{A}$ we have
$W_{\omega_{T},C}^{\mathrm{e}}(A,B;x)\equiv\omega_{T}^{\mathrm{e}}(A\tau_{x}(B))-\omega^{\mathrm{e}}(A)\omega^{\mathrm{e}}(B)\longrightarrow 0\text{ \qquad as\qquad\ }\left|x\right|\longrightarrow\infty$ (1.1.11)
The capital letter $W$ above refers to “Wightman” in order to emphasize the close connection with quantum field theory, see Subsection 1.4. Other terminologies in common use are “cumulants”, “truncated correlatin functions”, and “Ursell functions”. Equation (1.1.11) is a general property of a special class of states, namely those ergodic states $\omega^{\mathrm{e}}$, which are also in the subset of the set of extremal states which are invariant under a group (see, e.g. [Rue69], Definition 6.3.1). In our case, the group is that of lattice translations on $\mathbf{Z}^{\nu}$, and the extremal states are identified with the pure phases [Wight79]: they cannot be written as a convex combination of two other states $\omega_{1}\neq\omega^{\mathrm{e}}$ and $\omega_{2}\neq\omega^{\mathrm{e}}$, i.e., the relation $\omega^{\mathrm{e}}=a\omega_{1}+(1-a)\omega_{2}\text{ with }0<a<1$ does not hold unless $\omega_{1}=\omega^{\mathrm{e}}$ and $\omega_{2}=\omega^{\mathrm{e}}$. We now restrict our discussion to the phase-transition region of the Ising model, the extremal ergodic states being $\omega_{1}^{\mathrm{e}}=\omega_{T}^{+}$ and $\omega_{2}^{\mathrm{e}}=\omega_{T}^{-}$; these states may be obtained as limits (as previously described) of the states $\omega_{T}^{\Lambda}$, as $\Lambda\nearrow\mathbb{Z}^{\nu}$, with $+$ or -- b.c., respectively, and are not invariant under the spin-flip symmetry, i.e., as stated before, there is s.s.b.. If $0\leq T<T_{c}$, the closed segment
$\omega_{T}=a\omega_{T}^{+}+(1-a)\omega_{T}^{-}\qquad\text{ with }\qquad\ 0\leq a\leq 1$ (1.1.12)
describes, geometrically, the two-phase region, with boundary points the extremal ergodic states (pure phases) - in the lattice-gas language, the liquid phase and the vapour phase. The critical state $\omega_{T_{c}}$ is spin-flip invariant and unique, hence also ergodic - it may be obtained as the limit of the states $\omega_{T_{c}}^{\Lambda}$, with free or periodic b.c. on $\Lambda$.
Remark 1.1.
It should be emphasized that the concepts of extremal and ergodic states may be distinct, as, for instance, in the Ising antiferromagnet, for which there are two invariance groups, corresponding to the translations within each sublattice. In this paper we shall, however, restrict ourselves to the Ising universality class, for which the two concepts coincide, and use the words “extremal states” and “ergodic states” interchangeably.
We now come back to the precise definition of the dependence of the r.v. mentioned by Glimm. The probabilistic approach to the RG leads to conjecture that: a.) if $0<T<T_{c}$, the rate of clustering of the connected (Wightman) two-point functions for the ergodic states $\omega_{T,C}^{\pm}$ given by (1.1.11) is such that
$W_{\omega_{T}^{\pm},C}(A,B;x)\in l^{1}(\mathbb{Z}^{\nu})\qquad\text{ if }\qquad 0\leq T<T_{c}$ (1.1.13)
and b.) if $T=T_{c}$, i.e., at the critical point, the clustering is not summable:
$W_{\omega_{T_{c}},C}(A,B;x)\notin l^{1}(\mathbb{Z}^{\nu})$ (1.1.14)
Thus, the critical point is singled out among the ergodic states by the condition of non-summable clustering (1.1.14). In the one-dimensional $1/r^{2}$ model Imbrie and Newman ([IN88], see also [DHUM92]) have proved, however, that there is an intermediate regime $T_{0}<T<T_{c}$ with positive magnetization but nonsummable connected two-point functions.
Remark 1.2.
Although not for Ising models, it is possible that there is a first-order transition, as happens for n.n. high-q Potts models. At the “critical” temperature separating the low-T and high-T regime, there coexist q extremal (ergodic) ordered states with one ergodic (extremal) disordered state, all $q+1$ states with exponentially decaying correlations. We thank the referee for this remark.
We define the (finite region) susceptibility
$\chi_{\Lambda}(T)=\frac{\partial^{2}f_{\Lambda}}{\partial h^{2}}(T,0)$ (1.1.15)
By (1.1.3), this is a normalized fluctuation of the r.v. $\sigma_{x}^{3}$ , i.e.,
$\chi_{\Lambda}(T)=\omega_{T}^{\Lambda}\left(\left(\left.\sum\nolimits_{x\in\Lambda}\left(\sigma_{x}^{3}-\omega_{T}^{\Lambda}(\sigma_{x}^{3})\right)\right/|\Lambda|^{1/2}\right)^{2}\right)$ (1.1.16)
Near the critical point we have [Thomp72]
$\chi(T)\equiv\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\chi_{\Lambda}(T)\approx(T_{c}-T)^{-\gamma}\qquad\text{ as}\qquad T\nearrow T_{c}$ (1.1.17)
Equation (1.1.17) defines the critical exponent $\gamma$; at the same time, the spontaneous magnetization (1.1.7) has the behavior
$m_{s}\approx(T_{c}-T)^{\beta}\qquad\text{ as}\qquad T\nearrow T_{c}$ (1.1.18)
Equation (1.1.18) defines the critical exponent $\beta$. Equations (1.1.17) and (1.1.18) lead us to define the normalized fluctuation operator associated to the r.v. (operator) $\sigma^{3}$ by the equation
$\tilde{F}_{\Lambda,T}(\sigma^{3})\equiv\frac{1}{(\tilde{V}_{T}^{\Lambda})^{1/2}}\sum_{x\in\Lambda}(\sigma_{x}^{3}-\omega_{\Lambda}^{T}(\sigma_{x}^{3}))$ (1.1.19)
where
$\tilde{V}_{T}^{\Lambda}\equiv\omega_{T}^{\Lambda}\left(\left(\sum\nolimits_{x\in\Lambda}\left(\sigma_{x}^{3}-\omega_{T}^{\Lambda}(\sigma_{x}^{3})\right)\right)^{2}\right)$ (1.1.20)
is the variance of the “block-variable” $\sum_{x\in\Lambda}\sigma_{x}^{3}$ in the state $\omega_{\Lambda}^{T}$. For $0<T<T_{c}$, this is consistent with (1.1.13) and the normalization by $\sqrt{\left|\Lambda\right|})$ in (1.1.16), but, for $T=T_{c}$, one expects  anomalous fluctuations, defined by:
$V_{\Lambda}^{T}=O(|\Lambda|^{\alpha})\text{ \qquad with\qquad\ }\alpha>\frac{1}{2}\text{ \quad as \quad\ }\left|\Lambda\right|\longrightarrow\infty$ (1.1.21)
These anomalous fluctuations are associated to strong density fluctuations, whereby light is strongly scattered at the critical point: this is the beautiful phenomenon of critical opalescence, of which one of the most detailed theoretical discussions is Einstein’s early paper [Ein1909]. The Ising model, in its lattice-gas version for liquids, is the foremost model in this connection ([PV02], p. 552), providing ample empirical support to (1.1.21) [Fis67].
A different definition of normalized fluctuation operator is
$F_{\Lambda,T}(\sigma^{3})\equiv\frac{1}{\left(V_{\Lambda}^{T}\right)^{1/2}}\sum_{x\in\Lambda}(\sigma_{x}^{3}-\omega_{T}(\sigma_{0}^{3}))$ (1.1.22)
where
$V_{\Lambda}^{T}\equiv\omega_{T}\left(\left(\sum\nolimits_{x\in\Lambda}\left(\sigma_{x}^{3}-\omega_{T}(\sigma_{0}^{3})\right)\right)^{2}\right)$ (1.1.23)
Associated to $F_{\Lambda,T}(\sigma^{3})$ and $\tilde{F}_{\Lambda,T}(\sigma^{3})$ we have two types of characteristic functions $\Phi_{\Lambda,T},\tilde{\Phi}_{\Lambda,T}$, defined by (for $\omega_{T}$ an ergodic state):
$\Phi_{\Lambda,T}(t)\equiv\omega_{T}(\exp(itF_{\Lambda,T}))\ ,\qquad\forall t\in\mathbb{R}$ (1.1.24)
and
$\tilde{\Phi}_{\Lambda,T}(t)\equiv\omega_{T}^{\Lambda}((\exp(it\tilde{F}_{\Lambda,T}))\ ,\qquad\forall t\in\mathbb{R}$ (1.1.25)
The limits (for $\omega_{T}$ an ergodic state)
$\Phi_{T}(t)\equiv\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\Phi_{\Lambda,T}(t)\ ,\qquad\forall t\in\mathbb{R}$ (1.1.26)
and
$\tilde{\Phi}_{T}(t)\equiv\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\tilde{\Phi}_{\Lambda,T}(t)\ ,\qquad\forall t\in\mathbb{R}$ (1.1.27)
(provided they exist) need not coincide, see Subsection 1.3.3. Moreover, for instance, the limit (1.1.26) need not exist in general: for instance, taking random b.c. on $\Lambda$ on the r.h.s. of (1.1.26), for $T<T_{c}$, one may have several distinct subsequential limits along sequences of finite regions $\Lambda_{n},n=1.2,...$ of the r.h.s. of (1.1.26), see [vES02].
We now finish the summary of contents of our paper.
Since we shall base most of our discussion on the states of infinite systems, which are the generalizations of (1.1.10) to the whole quasi-local algebra $\mathcal{A}$, we make an interlude, in Subsection 1.2, in order to establish some of the fundamental notions.
Subsection 1.3 is concerned with phase transitions and spontaneous symmetry breaking (s.s.b.) in ferromagnetic spin systems. It is divided into four subsections.
Subsection 1.3.1 deals with generalities. Subsection 1.3.2 is concerned with the statement and comments on a criterion of criticality (Criterion A) based on (1.1.14). Subsection 1.3.3 defines SLRO, LLRO and discusses their connection with Criterion A. Subsection 1.3.4 introduces the connected (or truncated) $r$- point Wightman functions (also called Ursell functions) and some of the associated inequalities which are used in the text. Subsection 1.3.5 defines the concepts of interacting and quasi-free theories.
It is only in Subsection 1.4 that the main subject of this paper, that of the relation between macroscopic phenomena and the laws of microphysics, is introduced. It defines the fluctuation state $\omega_{T}^{F}$, the “state induced by the fluctuations” from an equilibrium state at temperature $T$, together with our main related assumptions, and its presumed connection with renormalization-group (RG) transformations. The possible extension of $\omega_{T}^{F}$ to the general “quantum Ising systems”, which constitute the “Ising universality class” associated to the Ising model, defined in Subsection 1.2, and the associated “Lie algebra of fluctuations” ([Ve11], [VZ92], [HL73], [Sew86]) is also briefly discussed there.
In Subsection 1.5 we study the ferromagnetic Ising model in finite volume. It is shown, as direct corollaries of the work of De Coninck and Gottal [dCG84] that the characteristic function $\Phi_{\Lambda,T}(t)$ defines an interacting system (equivalently: is Non-Gaussian) for all temperatures $0\leq T<\infty$. One further result in this framework, Proposition 1.6, is used in the proof of Theorem 2.6.
Section 2 is devoted to the connected Wightman functions ($r\geq 2$), both for $0<T<T_{c}$ and for $T=T_{c}$. Subsection 2.1 describes the framework which was introduced by one of us (M.R.) in [MR02] and reviews the case $0<T<T_{c}$, both for $r=2$ and $r>2$. In Subsection 2.2 we restrict ourselves to $r=2$ and $T=T_{c}$, establishing the existence of anomalous fluctuations for $\nu=2$ and for $\nu=3$, under the standard assumption (2.3.24).
In Subsection 2.3 we approach the question of the connected $r$-point function for $r\geq 4$ in the critical case. Under an assumption (Scaling Assumption C), which is the natural analogue of (2.3.16) for $r\geq 4$, we prove in Theorem 2.6 that the fluctuation state $\omega_{F}^{T}$ is quasi-free (corresponding to a Gaussian characteristic function). This conjecture seems to be widespread among several mathematical physicists (private communications by S. Molchanov and G. Slade to one of us (D.H.U.M.)), but the sole theorem we were able to found which claims to prove it was Theorem 6.2 in the paper by Cassandro and Jona-Lasinio [CasJL76]. Their proof is, however, incomplete (see also  1.7 and  1.8 ,in particular concerning the possible connection with one of the peculiarities pointed out by Griffiths and Pearce, and the important work of Wehrl).
In Section 3 we study several aspects of universality and its violation. Using the Dyson model as a prototype the known universality of critical exponents such as $\beta$ and $\gamma$ for long-range interactions, is briefly reviewed in Subsection 3.1. In Subsection 3.2 we argue that certain properties of energetic and dynamical stability, which distinguish long-range interactions from short-range ones, are among those which violate universality. In order to prove this conjecture in an explicit model (the Dyson model), we show that the Cloitre function, recently introduced by Albert and Kiessling ([AlKi17], see also [Ki13]), allows an exact calculation of “dynamical critical exponents”, which are seen not to be universal, and, moreover, are proved to be totally unaccounted for by (rigorous) nonequilibrium mean-field theory.
Section 4 is a conclusion, with a discussion of the main open problems.
Appendix A is devoted to an elementary proof [vHBW84] of the existence of a gap in the physical Hamiltonian associated to the forthcoming “Ising universality class” in Subsection 1.2, see also [KN97], [KN98], [KNS01], [NSS07], [DVEPR23], as well as the comments in Subsection 1.4 and Remark A.1 in Appendix A.
Appendix B provides (we hope) a comprehensive review of the beautiful work of Newman [Ne79], which is complementary to ours, in that it concerns the upper critical dimension.
Since we shall base our discussion on the states of infinite systems, which are the generalization of (1.1.10) to the whole “quasilocal” algebra $\mathcal{A}$, we need to make an interlude in the next subsection in order to establish some of the fundamental notions. The notions need not be very extended, because we shall restrict ourselves to the Ising universality class (the forthcoming generalized Ising model in the case of nearest-neighbor interactions, the Dyson model in the case of long-range interactions).
1.2 Brief mathematical interlude on states of infinite systems
In this section we briefly summarize some of the notions we shall use. The classic book [Sew86] is a comprehensive, elegant and deep survey. See also [Hug72], [LvH78], and the treatises [BRo297], [BRo186].
A prototype is the generalized Heisenberg Hamiltonian (generalized Ising model (gIm) if $\delta=0$):
$H_{\Lambda}=-2\sum_{x,y\in\Lambda}\left(j(\left|x-y\right|)(\sigma_{x}^{3}\sigma_{y}^{3}+\delta H_{XY}(x,y))\right)$ (1.2.1)
where
$H_{XY}(x,y)\equiv\sigma_{x}^{1}\sigma_{y}^{1}+\sigma_{x}^{2}\sigma_{y}^{2}$ (1.2.2)
and
$\delta\in(-1,1)$ (1.2.3)
and
$\sum_{x\in\mathbb{Z}^{\nu}}\left|j(x)\right|<\infty\text{ \qquad and\qquad\ }j(0)=0$ (1.2.4)
Above, $\left|x\right|\equiv\left(\sum_{i=1}^{\nu}x_{i}^{2}\right)^{1/2}$, $\nu$ denotes the dimension, and, as asserted in the introduction, the external field will be set equal to zero in the present paper. Equation (1.2.4) is the important requirement of thermodynamic stability, necessary for the existence of the thermodynamic limit of the energy per unit volume. Above, $\delta\in(-1,1)$ and we assume
$j(\left|x\right|)\geq 0$ (1.2.5)
which is appropriate to describe ferromagnetism. Above, too, $\sigma_{x}^{i}$, $i=1,2,3$, are the Pauli matrices at the site $x$, and $H_{\Lambda}$ acts on the Hilbert space ${\cal H}_{\Lambda}=\otimes_{x\in\Lambda}\mathbb{C}_{x}^{2}$, and $\vec{S}_{x}$ is short for $\mathbf{\ 1}\otimes\cdots\otimes\vec{S}_{x}\otimes\cdots\otimes\mathbf{1}$. We define the algebra of (bounded) operators acting on the Hilbert space ${\cal H}_{\Lambda}$: ${\cal A}(\Lambda)=B({\cal H}_{\Lambda})$. These local algebras satisfy: 1.) Causality: $[{\cal A}(B),{\cal A}(C)]=0$ if $B\cap C=\phi$ and 2.) Isotony: $B\subset C\Rightarrow{\cal A}(B)\subset{\cal A}(C)$. ${\cal A}_{L}=\cup_{B}{\cal A}(B)$ is termed the local algebra; its closure with respect to the norm, the quasilocal algebra (observables which are, to arbitrary accuracy, approximated by observables attached to a finite region). The norm is $B({\cal H}_{\Lambda})\ni A\longmapsto\left\|A\right\|=\sup_{\left\|\Psi\right\|\leq 1}\left\|A\Psi\right\|$, $\Psi\in{\cal H}_{\Lambda}$.
For a finite quantum spin system let $\rho_{\Lambda}$ denote a density matrix on ${\cal H}_{\Lambda}$, i.e., a non-negative Hermitian matrix of unit trace on ${\cal H}_{\Lambda}$. We may view $\rho_{\Lambda}$ as a state $\omega_{\Lambda}$ on $\cal{A}(\Lambda)$ – a positive, normed linear functional on ${\cal A}(\Lambda)$: $\omega_{\Lambda}(A)=\text{tr}_{{\cal H}_{\Lambda}}(\rho_{\Lambda}A)\text{ for }A\in{\cal A}(\Lambda)$ (positive means $\omega_{\Lambda}(A^{{\dagger}}A)\geq 0$, normed $\omega_{\Lambda}(\mathbf{1})=1$.)
This notion of state generalizes to systems with infinite number of degrees of freedom $\omega(A)=\lim_{\Lambda\nearrow\infty}\omega_{\Lambda}(A)$, at first for $A\in{\cal A}_{L}$ and then to ${\cal A}$.
Above, $\lim_{\Lambda\nearrow\infty}$ denotes the thermodynamic limit, in the simplest case along a sequence of parallelepipeds whose sides tend to infinity. We shall always consider that $\Lambda=-\Lambda\subset\mathbb{Z}^{\nu}$.
Considering quantum spin systems on $\mathbb{Z}^{\nu}$, we shall consider only space-translation-invariant states, i.e., such that
$\omega\circ\tau_{x}=\omega\qquad\text{ for all }\qquad x\in\mathbb{Z}^{\nu}$ (1.2.6)
As remarked in the previous subsection, an extremal invariant or ergodic state (see, however, Remark 1.1) is a state which cannot be written as a proper convex combination of two distinct states $\omega_{1}$ and $\omega_{2}$, i.e., the following does not hold:
$\omega=a\omega_{1}+(1-a)\omega_{2}\qquad\text{ with }\qquad 0<a<1$ (1.2.7)
If the above formula is true, it is natural to regard $\omega$ as a mixture of two pure “phases” $\omega_{1}$ and $\omega_{2}$, with proportions $a$ and $1-a$, respectively [Wight79]. Alternatively, (see [LvH78], p.403 and references given there), a state $\omega$ is ergodic with respect to space translations $\tau_{x}$ iff
$\lim_{\Lambda\nearrow\infty}\frac{1}{|\Lambda|}\sum_{x\in\Lambda}\omega(\tau_{x}(A)B)=\omega(A)\omega(B)$ (1.2.8)
The strong limit s-lim below is the limit when applied to any vector in the representation space. The following theorem is fundamental ([KaRo66], see also [LvH78] for an elementary proof):
Theorem 1.1.
Let $\omega$ be any translation-invariant state. Then the space average of A
$\eta_{\omega}(A)\equiv\operatorname*{s-lim}_{\Lambda\nearrow\infty}\frac{1}{|\Lambda|}\sum_{x\in\Lambda}\Pi_{\omega}(\tau_{x}(A))$ (1.2.9)
exists, and, if $\omega$ is ergodic, then
$\eta_{\omega}(A)=\omega(A)\mathbf{1}$ (1.2.10)
Equation (1.2.10) corresponds to “freezing” the observables at infinity to their expectation values. The space averages $\eta$ defined above correspond to macroscopic “pointer positions”, e.g., the mean magnetization in the Heisenberg model (1.2.1) in the $3$- direction
$m_{\Lambda}\equiv\frac{1}{|\Lambda|}\sum_{x\in\Lambda}\sigma_{x}^{3}$ (1.2.11)
with $A=\sigma^{3}$. If $\eta_{\omega_{+}}(\sigma^{3})=a_{+}>0$, and $\eta_{\omega_{-}}(\sigma^{3})=a_{-}<0$, the states $\omega_{\pm}$ are macroscopically different, i.e., differ from one another by flipping an infinite number of spins. For a comprehensive discussion, see [Sew86], Section 2.3.
Finally, a word about equilibrium states is necessary (for much more, see [Hug72]). There is an algebra $\tilde{{\cal A}}$ of analytic elements which is dense in ${\cal A}$, i.e., such that the function $\tau_{z}(A)$ is an entire function of the complex variable $z$ with values in ${\cal A}$, and reduces to the time-translation automorphism of ${\cal A}$ when $z$ is real (Lemmas 4.7 and 4.8 of [Hug72]). We say that a state $\omega$ is a thermal equilibrium state of the system at inverse temperature $\beta$, or, equivalently, a KMS state with respect to $\tau_{t}$ if
$\omega(\tau_{t}(A)B)=\omega(B\tau_{t+i\beta}(A))\qquad\forall B\in{\cal A}\text{ \quad and \quad}A\in\tilde{{\cal A}}$ (1.2.12)
The Gibbs state of a finite system at inverse temperature $\beta$ is given by
$\omega_{\beta,\Lambda}(A)=\text{tr}(\rho_{\Lambda}A)\qquad\text{ with \quad}\rho_{\Lambda}=\text{ const.}\exp(-\beta H_{\Lambda})$ (1.2.13)
It is immediate that (1.2.13) satisfies (1.2.12), but (1.2.12) also characterizes thermal equilibrium states of infinite systems, which will be most frequently (but not exclusively, see Subsection 3.2) considered in the present paper.
1.3 Phase transitions in ferromagnetic lattice systems
1.3.1 Generalities
For the model given by (1.2.1), (1.2.4) and (1.2.5) of ferromagnetic lattice systems, we shall say that the interactions are  short-range if the potential $j$ decays exponentially, i.e.,
$\left|j(\left|x\right|)\right|\leq c\exp(-\left|x\right|/\Gamma)$ (1.3.1)
for some $c,\Gamma$ both positive constants. A good approximation of (1.3.1) is a finite-range interaction, i.e., $j(\left|x\right|)=0$ for some $\left|x\right|\geq e$, $e$ being a fixed real number, or even a nearest-neighbor interaction
$j(\left|x\right|)=0\text{ \qquad unless }\qquad\left|x_{i}\right|=1\text{ for all }i=1,\ldots,\nu$ (1.3.2)
We shall say that the interactions are long-range if they are not finite-range and there is polynomial decay, i.e.,
$j(\left|x\right|)=O(\left|x\right|^{-\alpha})\text{ \qquad as\ }\left|x\right|\longrightarrow\infty\text{ \ for some \ }\alpha>\nu$ (1.3.3)
Equation (1.3.3) is due to the stability condition (1.2.4). We define the spin-reflection symmetry $\alpha$ by the following equation, valid for all $x\in\mathbb{Z}^{\nu}$:
$\alpha(\sigma_{x}^{1})=\sigma_{x}^{1}\text{ }\qquad\text{together with }\qquad\alpha(\sigma_{x}^{2})=-\sigma_{x}^{2}\text{ \ and\ }\alpha(\sigma_{x}^{3})=-\sigma_{x}^{3}$ (1.3.4)
We have the following theorem:
Theorem 1.2.
For the model defined by equations (1.2.1), (1.2.3), (1.2.4), (1.2.5) and (1.2.6), assume $\nu\geq 2$. Then, there exists a $T_{c}(\nu)>0$ such that there exist two translation-invariant (extremal) ergodic equilibrium (KMS) states $\omega_{\pm,T}^{e}$ for all $T<T_{c}(\nu)$ which are not invariant under the spin-flip symmetry (1.3.4), and we say that a spontaneous symmetry breakdown (s.s.b.) occurs. The mean magnetizations $m_{\pm,T}$ associated to these states are given by $\eta_{\omega_{\pm,T}^{\mathrm{e}}}(\sigma^{3})\equiv m_{\pm,T}$ with $m_{+,T}>0$ and $m_{-,T}=-m_{+,T}$.
Proof. The statements of the theorem follow from Theorem 6.2.48 of [BRo297], together with the remarks before that theorem, and Theorem (1.1). The assumed translation invariance of the states is important: see the remarks on the first paragraph of p.338 of [BRo297] and the references given there.
$\Box$
Remark 1.3.
The case of $\delta=0$ corresponds to the Ising model. We shall consider this boundary case in most of the paper, but follow ([BRo297], p.320 et seq.) in considering the associated states as states on the whole algebra $\mathcal{A}$ as before. This choice agrees with the physical significance of the model in applications, which relies on the fact that it is an anisotropic limit ($\delta\rightarrow 0$) of the model (1.2.1). The dynamics of Subsection 3.2 also takes this choice into account. When restricting to observables which are functionals of the $\sigma_{x}^{3}$; $x\in\mathbb{Z}^{\nu}$, however, the correlation inequalities which are specific to the Ising model, remain, of course, applicable, as well as the probabilistic language, commonly associated with the “classical version”.
Theorem 1.2 substantiates considering the model defined by equations (1.2.1) as belonging to the Ising universality class. An additional reason for this nomenclature is the property of Ising domination proved in Appendix A.
Although we keep the dimension $\nu$ general, for clarity, we shall only consider , when dealing with short-range interactions, the (physically relevant) cases $\nu=2$ and $\nu=3$ in correspondence with Theorem 1.2. An upper bound to the upper critical dimension is $\nu_{c}=4$ [ADC21]. For high dimensions $\eta=0$, and the systems rescales to a “massless” Gaussian, which is quasi-free, but has connected correlations decaying as $r^{2-\nu}$. The very recent paper by Duminil-Copin and Panis [DCP24] proves that the critical exponent $\eta=0$ for $\nu$ at least 4, and gives improved rigorous bounds on $\eta$ for $\nu=3$.
1.3.2 A criterion of criticality: Criterion A
The critical state $\omega_{T_{c}}$ is the unique, spin-flip invariant – and hence ergodic – state at $T=T_{c}$. By Example 4.3.24 of [BRo186], together with the states $\omega_{\pm,T}^{\mathrm{e}}$ in Theorem 1.2, it also satisfies the property of clustering in space:
$\lim_{\left|x\right|\rightarrow\infty}\omega_{T,C}^{\mathrm{e}}(\sigma^{3}\tau_{x}(\sigma^{3}))=0$ (1.3.5)
with the notation (1.1.11). The above equation (1.3.5) is therefore valid for all $0<T\leq T_{c}$. It is now natural to inquire into the rate of clustering. It has recently been proved in a remarkable paper [DCGR20] that
$0\leq\omega_{T,C}^{\mathrm{e}}(\sigma^{3}\tau_{x}(\sigma^{3}))\leq\exp(-c_{T}|x|)\qquad\forall x\in\mathbb{Z}^{\nu}$ (1.3.6)
with
$c_{T}>0\qquad\forall\quad 0<T<T_{c}\ .$ (1.3.7)
The quantity $c_{T}$ in (1.3.6) has the interpretation of “inverse correlation length”, and, since it is also strictly positive for $T>T_{c}$ (see [DCGR20] and references given there), it may tend to zero only at the critical point. This possibility, and its implications for the phenomenological theory of scaling and critical fluctuations (see [Sew86], 5.4.1, and references given there), will be one of our major concerns in the present paper: it has been analysed by Sewell [Sew86], p. 122, section 5.4), who suggested the following argument.
Let
$\bar{\omega}_{T}\equiv\frac{1}{2}(\omega_{+,T}^{\mathrm{e}}+\omega_{-,T}^{\mathrm{e}})$ (1.3.8)
The state $\bar{\omega}_{T}$ is spin-flip and translation invariant, and, by Theorem 1.2 it follows immediately that
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\frac{1}{|\Lambda|^{2}}\bar{\omega}_{T}\left(\left(\sum\nolimits_{x\in\Lambda}\sigma_{x}^{3}\right)^{2}\right)=\left|m_{\pm,T}\right|^{2}>0$ (1.3.9)
as long as $0<T<T_{c}$, and, hence,
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\frac{1}{\left|\Lambda\right|}\sum_{x,y\in\Lambda}\bar{\omega}_{T}(\sigma_{x}^{3}\sigma_{y}^{3})=\infty$ (1.3.10)
for $T<T_{c}$. By translation invariance of $\bar{\rho}$, (1.3.10) implies that
$\sum_{y\in\mathbb{Z}^{\nu}}\bar{\omega}_{T}(\sigma_{0}^{3}\sigma_{y}^{3})=\infty\text{ for }T<T_{c}$ (1.3.11)
Since $\bar{\omega}_{T}(\sigma_{x}^{3})=0$, $\forall x\in\mathbb{Z}^{\nu}$, it is natural to assume that (1.3.11) holds, by continuity, up to the critical point, the critical state being unique and thus also invariant under the spin-flip symmetry:
$\sum_{x\in\mathbb{Z}^{\nu}}\omega_{T_{c}}^{\mathrm{e}}(\sigma^{3}\tau_{x}(\sigma^{3}))=\infty$ (1.3.12)
i.e., the “clustering is not summable at the critical point”. Since $\omega_{T_{c}}$, being unique, is also ergodic, equations (1.3.6), (1.3.7) and (1.3.12) suggest the following criterion:
Criterion A The critical state is singled out, among all ergodic states for $T\leq T_{c}$, by the condition that its connected two-point function is not summable.
Criterion A is not new: it agrees with the “renormalization group and probability theory” approach of several researchers, see the review by Jona-lasinio [JL01] and references given there.
Remark 1.4.
The uniqueness of critical state is related with another important issue regarding the continuity of spontaneous magnetization for ferromagnetic Ising model, established recently for $d=3$. See [ADCV15] and references therein.
1.3.3 SLRO, LLRO and their connection with Criterion A
We now come back to the notions of normalized fluctuation operator (1.1.19), (1.1.20), (1.1.22) and (1.1.23), and the corresponding characteristic functions (ch.f.) $\Phi_{\Lambda,T}(t)$ and $\tilde{\Phi}_{\Lambda,T}(t)$, given by (1.1.24) and (1.1.25), respectively. The two different limits $\Phi_{T}$ and $\tilde{\Phi}_{T}$, given by (1.1.26) and (1.1.27) (when they exist) are known as short-long-range-order (SLRO) and long-long-range order (LLRO), respectively [SML64]. They need not coincide - the fact that the normalized fluctuation operators are not uniformly bounded in norm leaves this as an a priori possibility, but explicit examples have in fact been given in mean-field theories, by Ellis, Newman and Rosen [ENR80], see also the explicit example in [CasJL78], which has been abstracted from [ENR80]. We conjecture, however, that LLRO and SLRO agree for short-range interactions. This conjecture is relevant to Theorem 2.6.
Remark 1.5.
In order to analyse the transition from $0<T<T_{c}$ to $T=T_{c}$, which is implicit in Criterion A, it is necessary (see [DCGR20]) to consider one of the ergodic states $\omega_{\pm,T}^{e}$ (for which the mean magnetization $m\neq 0$), and $\omega_{T_{c}}$, for $T=T_{c}$. If the alternative $\Phi_{\Lambda,T}$, as $\Lambda\nearrow\mathbb{Z}^{\nu}$, is considered (equivalent to LLRO [SML64]), boundary conditions $\pm$ must be taken at $\Lambda$ for $T<T_{c}$, and free or periodic b.c. at $T=T_{c}$, and a fixed choice, which imposes itself to define the procedure in precise mathematical terms is not adequate to cover both cases. Since different ergodic states of infinite systems are typically associated to performing the thermodynamic limit with different choices of b.c., as indicated above, we see that, in addition to the arguments employed in Subsection 1.1, the choice of regarding the states of infinite systems as the basic objects of the theory is the only natural one.
1.3.4 The connected Wightman functions
We now come back to (1.1.24) and (1.1.26) and define
$\left.\frac{d^{r}}{dt^{r}}\log\Phi_{A,\Lambda}(t)\right|_{t=0}=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{r}:\\ x_{i}\in\Lambda,~{}i=1,...,r\end{subarray}}W_{\omega_{T},C}^{r}\left(x_{1},\ldots,x_{r}\right)$ (1.3.13)
where the $W_{\omega_{T},C}^{r}(x_{1},\cdots,x_{r})$ are the connected (or truncated) Wightman (or correlation) $r$-point functions associated to the state $\omega_{T}$:
$W_{\omega,C}^{1}\left(x_{1}\right)\equiv\omega\left(A_{x_{1}}\right)$ (1.3.14)
$W_{\omega,C}^{2}\left(x_{1},x_{2}\right)\equiv\omega\left(A_{x_{1}}A_{x_{2}}\right)-\omega\left(A_{x_{1}}\right)\omega\left(A_{x_{2}}\right)$ (1.3.15)
and, generally, by recursion,
$W_{\omega,C}^{r}\left(x_{1},\ldots,x_{r}\right)\equiv\omega\left(A_{x_{1}}\cdots A_{x_{r}}\right)-\sum_{\pi_{1},\ldots,\pi_{k}}\prod_{j=1}^{k}W_{\omega,C}^{j}\left(x_{i_{1}}\ldots x_{i_{j}}\right)$ (1.3.16)
at $\omega=\omega_{T}$, where the sum in (1.3.16) runs over all non-trivial (i.e., $k>1$) partitions of $x_{1},\ldots,x_{r}$ into subsets $\pi_{1},\ldots,\pi_{k}$. We replaced the notation $\Phi_{\Lambda}(t)$ in (1.1.24) by $\Phi_{A,\Lambda}(t)$, specifying the observable $A$ which is arbitrary, but in (1.1.24) was $A=\sigma^{3}$. This choice will be made in this section, because we shall consider the Ising model ($\delta=0$ in (1.2.1)), but keeping it general at this stage serve the purpose of clarity. Notice that (1.3.15) is consistent with (1.1.11), in this notation, with $A=B$.
If $A=\sigma^{3}$, and $\delta=0$ in (1.2.1), $W_{\omega_{T},C}^{r}$ where $\omega_{T}$ is any equilibrium state at temperature $T$ of the Ising model, two very important inequalities will concern us:
$W_{\omega,C}^{2}(x_{1},x_{2})\geq 0$ (1.3.17)
and
$W_{\omega,C}^{4}(x_{1},\ldots,x_{4})\leq 0$ (1.3.18)
equation (1.3.17) is Griffiths’ second inequality [Gri67], equation (1.3.18) is Lebowitz’ inequality [Leb74]. See also [GJ81].
Shlosman [Shl86] proved the general statement about the signs for higher-order Ursell functions.
1.3.5 Interacting and quasi-free theories
Some of the most difficult problems in the theory of the critical point concern the issue of triviality and the connections with quantum field theory. In the lattice-gas picture, model (1.2.1) has an interesting interpretation (see [BRo297], p.425 et seq): the XY term (which multiplies $\delta$ in (1.2.1)) is a discrete form of the Laplacean and is interpretable as a quasi-free or non-interacting lattice gas, while the Ising term corresponds to a potential energy arising from a two-body interaction between neighboring particles.
In the lattice-gas picture, with the spin-raising and lowering operators ([BRo297], p. 425) (in dimension $\nu=1$)
$a_{x}^{\ast}=\frac{\sigma_{x}^{1}+i\sigma_{x}^{2}}{2}$ (1.3.19)
and
$a_{x}=\frac{\sigma_{x}^{1}-i\sigma_{x}^{2}}{2}$ (1.3.20)
there arises a Hamiltonian term
$H=-2J\sum_{x=-n}^{n}a_{x}^{\ast}a_{x+1}^{\ast}a_{x+1}a_{x}$
which corresponds to a potential energy from a two-body interaction between neighboring particles, in contrast to $H_{XY}$ in (1.2.2), which is quadratic in the creation and annihilation operators $a_{x}$, $a_{x}^{\ast}$, $x\in\mathbb{Z}$, see ([BRo297], p.426). This latter case is an example of a quasi-free system, the former of an interacting system. We define precisely:
Definition 1.3.
A system defined by the connected Wightman functions $W_{\omega_{T},C}^{r}\left(x_{1},\ldots,x_{r}\right)$, given by (1.3.14), (1.3.15), and (1.3.16) is called quasi-free if $W_{\omega_{T},C}^{r}\left(x_{1},\ldots,x_{r}\right)\equiv 0$ for $r>2$, otherwise it is called an interacting system.
The thermal Wightman functions of the Ising model ((1.2.1) with $\delta=0$) are expected to define an interacting system in the sense of Definition  1.3, except if $T=0$. In this case, the ergodic states are (infinite) product states, see also Appendix A. Otherwise, the connected four-point Wightman function is not expected to be identically zero. This is because this would entail that all higher connected functions vanish, and hence already for finite $|\Lambda|$ the characteristic function would be a Gaussian, a behavior not expected for the Ising model which may be viewed as a system of interacting “fermions” in the sense of ([BRo297], p. 426).
Remark 1.6.
The importance of the interacting nature of the Ising system in our case is that the associated theory of phase transitions is physically a sound one, i.e., there are no non-physical features, as, e.g., in the phase transition in the free Bose gas, due to infinite compressibility, see, for instance, [Dor99], pp. 200-209. It is also related to the previously mentioned success of the Ising model in the description of real experiments. This property concerns the equilibrium states; in this paper, we shall be concerned, however, with the quasi-free or interacting nature of the macroscopic states, as explained in the next subsection.
1.4 Macroscopic phenomena
One of the deepest problems in physics is how macroscopic phenomena can be interpreted from the laws and structures of microphysics [Sew02]. One feature of this problem is to account for the observed macroscopic irreversibility, while the equations of microphysics are time-reversal invariant ([Wre20], [Wre23]). Another one is the construction of a sensible macroscopic measuring apparatus in the theory of quantum mechanical measurement (see [Wre123]: there a sequence of pure states for a system in finite volume tends to a mixed state in the thermodynamic limit.
Two observables reflecting macroscopic behavior are intensive observables, such as the mean magnetization (1.2.11), and the  fluctuation observables, such as (1.1.19) and (1.1.22).
A state which is the limit of
$\left(\left(a\Psi_{1,\Lambda}+b\Psi_{2,\Lambda}\right),(.)\left(a\Psi_{1,\Lambda}+b\Psi_{2,\Lambda}\right)\right)$
where $a$ and $b$ are complex numbers, such that
$|a|^{2}+|b|^{2}=1$
becomes, in the limit $\Lambda\nearrow\mathbb{Z}^{\nu}$, under certain conditions (see [Wre123] and references given there) a state of the form
$|a|^{2}(\Psi_{1},(.)\Psi_{1})+|b|^{2}(\Psi_{2},(.)\Psi_{2})$
That is, the relative phase of $b$ and $a$ in $a\Psi_{1,\Lambda}+b\Psi_{2,\Lambda}$ disappears in the (proper) macroscopic limit, leading to the mixed state (1.2.7).
In the theory of phase transitions (see Subsection 1.3.1, Theorem 1.2 shows that (1.2.11), the mean magnetization is an order parameter which distinguishes the two microscopic states $\omega_{\pm,T}^{\mathrm{e}}$ (phases) for $0<T<T_{c}$, and
$\omega_{T}\equiv a\omega_{+,T}^{\mathrm{e}}+(1-a)\omega_{-,T}^{\mathrm{e}}\text{ \qquad with \qquad}0<a<1$ (1.4.1)
represents a classical mixture. This remark applies to both the magnetic and the lattice fluid pictures, of course, but, in the latter case, when the temperature $T\nearrow T_{c}$, the resulting mixture displays an outstanding classical (macroscopic) effect - the critical opalescence - discussed in Subsection 1.1. In the former case, it is known that domains of a magnet are, in nature, found only in a definite direction. This constitutes one of the rare examples for which the existence of so-called superselection sectors [Wight95] originating from interactions with the environment, which are observed in nature, may be proved: a rigorous model of this situation may be constructed precisely around (1.2.1) following the model suggested by Narnhofer and Thirring [NT99], see [Wre23], Section 4.2, and references given there, as well as [NT99] and [NW14]. An alternative approach to this problem was suggested in [MR02], in which it was shown that every generic many-body system contains within the class of microscopic quantum observables a subalgebra of macro-observables, the spectrum of which comprises the macroscopic properties of the many-body system (see also [GL61]). A third, independent, approach to show the existence of domain walls was initiated in [ASW75] and further considerably extended by Koma, Nachtergaele, Spitzer and Starr ([KN97], [KN98], [KNS01], [NSS07]), see also Remark A.1 in Appendix A.
Concerning the fluctuations (1.1.19), (1.1.22), it is remarkable that they retain a quantum character in spite of the thermodynamic limit: in fact, they are related to (Bosonic) spin waves ([Ve11], [MR02]). consider the model (1.2.1) and, coming back to (1.3.19), (1.3.20), define
$a_{\Lambda}\equiv\frac{1}{\sqrt{\left|\Lambda\right|}}\sum_{x\in\Lambda}a_{x}$ (1.4.2)
and
$a_{\Lambda}^{\ast}\equiv\frac{1}{\sqrt{\left|\Lambda\right|}}\sum_{x\in\Lambda}a_{x}^{\ast}$ (1.4.3)
We have, by Theorem 1.1 and Theorem 1.2,
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\omega_{\pm,T}^{\mathrm{e}}([a_{\Lambda},a_{\Lambda}^{\ast}])=m_{\pm,T}$ (1.4.4)
since $[a_{\Lambda},a_{\Lambda}^{\ast}]=\left.\sum_{x\in\Lambda}\sigma_{x}^{3}\right/\left|\Lambda\right|$. Thus, for $0<T<T_{c}$, the quantities
$\left(\left|m_{\pm}\right|\left|\Lambda\right|\right)^{-1/2}\sum_{x\in\Lambda}\exp(ik\cdot x)a_{x}$ (1.4.5)
and
$\left(\left|m_{\pm}\right|\left|\Lambda\right|\right)^{-1/2}\sum_{x\in\Lambda}\exp(-ik\cdot x)a_{x}^{\ast}$ (1.4.6)
with $-\pi\leq k_{i}<\pi,i=1,\ldots,\nu$, are expected to define Boson (spin-wave) operators $a_{B}(k)$, $a_{B}(k)^{\ast}$ in the limit $\Lambda\nearrow\mathbb{Z}^{\nu}$, in the representation spaces of the states $\omega_{+,T}^{\mathrm{e}}$, for $0<T<T_{c}$, with $a_{B}^{\ast}$ replacing $a_{B}$ in the state $\omega_{-,T}^{\mathrm{e}}$. Equations (1.4.5), (1.4.6) exhibit the basic role of fluctuation operators as the links between quantum mechanics and macroscopic systems. This fact has been specially emphasized in [Ve11] and [MR02], and, previously, by Hepp and Lieb in their seminal paper on phase transitions out of thermal equilibrium [HL73]. There, they define precisely the limiting state $\omega_{T}^{F}$ (see also [MR02]):
Definition 1.4.
Assume that the limits
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\omega_{T}\left(F_{\Lambda,T}^{n}\right)\qquad\text{ with \qquad}n=1,2,3,\cdots$ (1.4.7)
exist. Above, if $\delta=0$ in (1.2.1), $F_{\Lambda,T}$ are defined by (1.1.22) if SLRO is chosen, otherwise (1.1.19) must be chosen, with $\omega_{T}$ replaced by $\omega_{\Lambda,T}$ in (1.4.7). We assume here the obvious extension of the definition to the full model (1.2.1), with three fluctuation operators $F_{\Lambda,T}^{i},i=1,2,3$ defined in analogy to (1.4.7):
$F_{\Lambda,T}^{i}\equiv\frac{1}{\sqrt{\left|\Lambda\right|}}\sum_{x\in\Lambda}\left(\sigma_{x}^{i}-\omega_{T}\left(\sigma_{0}^{i}\right)\right)\qquad\text{ with \quad}i=1,2,3$ (1.4.8)
Then, the state $\omega_{T}^{F}$, which we call the fluctuation state, is the state defined by the GNS construction (see [BRo186], p.153) on the polynomial algebra $\mathcal{P}$ [HL73], generated by new limiting objects, the fluctuation operators $F_{T}$, by
$\omega_{T}^{F}\left(\left(F_{T}^{i}\right)^{n}\right)=\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\omega_{T}\left(\left(F_{\Lambda,T}^{i}\right)^{n}\right)\qquad\text{ with \quad}n=1,2,\ldots\quad\text{and}\quad i=1,2\ .$ (1.4.9)
Note that the original equilibrium state $\omega_{T}$ is a state on a different algebra of observables, the quasi-local algebra $\mathcal{A}$.
We know of only two concrete constructions of fluctuation states $\omega_{T}^{F}$: Hepp and Lieb’s [HL73] and Wehrl’s [We71].
The spin-wave operators (1.4.5), (1.4.6) provide an accurate description of the low-temperature magnetization of ferromagnets [Mat65]. More recently, the temperature dependence of the junction critical Josephson current in the strong-coupling regime was obtained by using a quantum fluctuation description of charge qubits [BCFNV23], whose basic elements were collective quasi-spin fluctuation operators, which generate the Heisenberg algebra on the circle in a model of the strong-coupling BCS type (see, e.g., [Dor99], Chapter 35).
The convergence of (1.4.5), (1.4.6) to Boson spin-wave operators $a_{B}(k)$, $a_{B}^{\ast}(k)$ in the representation space of the states $\omega_{\pm,T}^{\mathrm{e}}$ is true under the assumption (1.4.7) of Definition 1.4 for, then, the limiting objects $F_{T}^{1,2}$ must satisfy (1.4.4). For $0\leq T<T_{c}$, the operators $F_{\Lambda,T}^{3}$ are defined as in (1.4.8), with the normalization by $1/\sqrt{\left|\Lambda\right|}$; as such, they define a Lie algebra, the Lie algebra of fluctuations ([Ve11], [VZ92]), whose importance was also emphasized in [HL73]; in the present case, it is the algebra $su(2)$, the Lie algebra of the group $SU(2)$. If we assume that, in addition to (1.4.4) and the fact that $m_{\pm,T}=0$ at $T=T_{c}$,
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\omega_{\pm,T}^{\mathrm{e}}(\left.\sigma_{x}^{i}\right/|\Lambda|)=0\qquad\text{ for \quad}i=1,2\,$ (1.4.10)
it follows that the thermodynamic limit induces a contraction of Lie algebra representations (see [CW99], [CW79]), viz. that of
$su(2)\rightarrow h(2)$ (1.4.11)
([CW79], III.1.2), $h(2)$ denoting the two-dimensional Heisenberg (CCR) algebra. Since $m_{\pm,T}=0$ at $T=T_{c}$, and assuming anomalous fluctuations (1.1.21) in the 3-direction, the passage to $T\nearrow T_{c}$ leads to an additional contraction
$h(2)\rightarrow r^{3}$ (1.4.12)
([CW79], III.1.3), where $r^{3}$ denotes the three-dimensional abelian Lie algebra, isomorphic to the Lie algebra of $\mathbb{R}^{3}$. If, in (1.4.8) the factors $1/\sqrt{\left|\Lambda\right|}$ are both replaced with $1/\left|\Lambda\right|^{\alpha}$, with $\alpha>1/2$, i.e., the fluctuations of $\sigma^{1,2}$ are also anomalous, (1.4.12) is also true a fortiori: see [Sew86], Appendix B, p.166, for a general theorem in this connection. Equations (1.4.11) and (1.4.12) together imply that the critical fluctuation state $\omega_{T}^{F}$ of the ferromagnets in the Ising universality class (1.2.1) is characterized, from the point of view of the algebra of fluctuations, as the unique state for which the Lie algebra of fluctuations is classical. This explains why critical behavior in quantum models may be described by classical Ising models.
We refer also to the references on the extensive work on the quantum central limit theorem given in the book by Verbeure [Ve11].
Remark 1.7.
A final, most important, aspect of the relation between the states $\omega_{T}^{F}$ and $\omega_{T}$ is the possible occurrence of a basic structural difference between them, taking into account, of course, that they are states on different algebras of observables, $\mathcal{P}$ and $\mathcal{A}$, as previously remarked. One possibility, proved under certain natural assumptions in the forthcoming Theorem 2.6, is that $\omega_{T}$ is an interacting state, while $\omega_{T}^{F}$ is a quasi-free state. We shall now try to provide some intuition, why this may be expected to occur.
The bosonic quasi-spin operators (1.4.5), (1.4.6) (spin-waves, with $\left|m_{\pm}\right|=1$) have been studied in great detail by Wehrl in the BCS model [We71]. In a certain limit (the Trotter limit, also used in [CW99], [CW79]), their dynamics is shown to satisfy the KMS condition (1.2.12) only if the value $k=0$ is excluded in (1.4.5), (1.4.6). The zero-momentum fluctuation operators are invariant under all permutations of the elements of the set $\{x\in\Lambda\}$, as is its Hamiltonian (which is the BCS mean-field Hamiltonian), and, comparing this situation with ordinary many-body systems, one sees that the operators $a_{\Lambda}$, $a_{\Lambda}^{\ast}$, in (1.4.2), (1.4.3), are the analogues of the center of mass variables, and thus behave freely. Consequently, they do not “thermalize”, and this happens for all temperatures. If we believe that these considerations apply, qualitatively speaking, to general models, we are led to conjecture that, for all temperatures (including $T_{c}$), the fluctuation state is quasi-free.
Furthermore, if we assume that the fluctuation state $\omega_{T}^{F}$ is a rigorous version of a “renormalization-group (RG) transformation” (in real space), the last assertion in the previous paragraph implies that RG transformations are singular, in the sense that $\omega_{T}^{F}$ and $\omega_{T}$ display a basic structural difference. Moreover, not being an equilibrium (KMS) state for any temperature (since any such state is an interacting state), it need not be locally Gibbs ((1.2.13)). The remarks in this paragraph provide a (possible) connection to the RG peculiarities pointed out by Griffiths and Pearce [GP79], see also van Enter’s comprehensive review [vE99] and the huge literature cited there, and Sewell’s pedagogic account in ([Sew86], 5.4.4).
1.5 Ferromagnetic Ising systems in finite volume and infinitely divisible characteristic functions
In this section we study the characteristic function $\tilde{\Phi}_{\Lambda,T}(t)$, given by (1.1.25), for finite volume. Our basic reference is the important paper of De Coninck and Gottal [dCG84]. Recall that a characteristic function (ch.f.) $f(t)$ is a Fourier-Stieltjes transform of a distribution function $dF$, defined for all $t\in\mathbb{R}$ by ([Chu74], Chap. 6.1, p. 142)
[…]
A ch.f. is called infinitely divisible iff, for all integers $n\geq 1$, there exist ch.f. $f_{n}$ such that
$f(t)=\left(f_{n}(t)\right)^{n}$ (1.5.2)
(see [Chu74], 7.6, or [Lu70], 108, 109). Instead of $\tilde{\Phi}_{\Lambda,T}(t)$, we consider
$\Phi_{T}^{\Lambda}(t)\equiv\exp\left(it\sum_{x\in\Lambda}\omega_{\Lambda,T}\left(\sigma_{x}^{3}\right)\right)\tilde{\Phi}_{\Lambda,T}(t)\,,$ (1.5.3)
or
$\Phi_{T}^{\Lambda}(t)=\omega_{\Lambda,T}\left(\exp\left(it\sum\nolimits_{x\in\Lambda}\sigma_{x}^{3}\right)\right)\ .$ (1.5.4)
For finite $\Lambda$, and real $t$, we have ([dCG84], Lemma 1):
$\Phi_{\Lambda}^{T}(t)=\prod_{j=1}^{\infty}\left(1+\frac{t^{2}}{t_{j}^{2}}\right)$ (1.5.5)
where $\left\{t_{j}\right\}_{j={1}}^{\infty}$ is an infinite set of real numbers such that
$0<t_{1}\leq t_{2}\leq\cdots$ (1.5.6)
Notice that (1.5.5) is invariant by inclusion of the variance in (1.1.19), since $t\rightarrow t/\sqrt{V_{\Lambda,T}}$ together with $t_{j}\rightarrow t_{j}/\sqrt{V_{\Lambda,T}}$. The following arguments are due to De Coninck and Gottal in [dCG84]. By (1.5.5), the reciprocal of $\Phi_{\Lambda}(t)$,
$\frac{1}{\Phi_{T}^{\Lambda}(t)}=\prod_{j=1}^{\infty}\left(1+\frac{t^{2}}{t_{j}^{2}}\right)^{-1}$ (1.5.7)
De Coninck and Gottal noticed that the quantity $\left(1+t^{2}/t_{j}^{2}\right)^{-1}$ is the ch.f. of the Laplace probability distribution, viz. $\exp(-\left|x\right|)$, over $x\in\mathbb{R}$, which is infinitely divisible ([Lu70], pp. 108, 109). Since the limit of a sequence of infinitely divisible ch.f. is infinitely divisible ([Lu70], Theorem 5.3.3, p. 110), $1/\Phi_{\Lambda}^{T}(t)$ is infinitely divisible. This result is the essential ingredient of Theorem 2 of [dCG84], whose consequences will be mentioned shortly. By (1.5.4) and (1.3.13), we have, connecting our notation with that ($U_{r,\Lambda}(0)$) of [dCG84],
$\left(\frac{d^{r}}{dt^{r}}\log\Phi_{\Lambda}^{T}\right)(0)$ $=$ $-U_{r,\Lambda}(0)$
$\equiv$ $i^{r}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{r};\\ x_{i}\in\Lambda,~{}i=1,...,r\end{subarray}}W_{\omega_{\Lambda,T},C}^{r}\left(x_{1},\ldots,x_{r}\right)$
Correspondingly,
$U_{r,\Lambda}(0)=-i^{r}\sum_{x_{1},\ldots,x_{r};x_{j}\in\Lambda}W_{\omega_{\Lambda,T},C}^{r}\left(x_{1},\ldots,x_{r}\right)$ (1.5.8)
Notice the difference of sign with respect to the previous equation, which corresponds to taking the reciprocal in (1.5.7). By Theorem 2 of [dCG84],
[…]
where
$K_{\Lambda}(+\infty)=2\sum_{j=1}^{\infty}\frac{1}{t_{j}^{2}}$ (1.5.10)
We have thus the important corollary
Proposition 1.5.
$\left|U_{2r+2,\Lambda}(0)\right|\neq 0$ for all $r=0,1,2,\cdots$
Proof. By (1.5.9), $U_{2r+2,\Lambda}\equiv 0$ iff $dK_{\Lambda}(x)=\lambda\delta(x)$ for some $\lambda>0$, and $\delta$ the delta measure at zero, but $K_{\Lambda}(\infty)\neq 0$ by (1.5.10).
$\Box$
A second result which will be useful for the forthcoming Theorem 2.6 is:
Proposition 1.6.
Assume that:
[…]
Then, if
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}U_{2r,\Lambda}(0)\equiv 0$ (1.5.12)
it follows that
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}U_{2(r-1),\Lambda}(0)\equiv 0\quad\text{for all }\quad\ r=3,4,5,\cdots$ (1.5.13)
and, consequently,
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}U_{2r,\Lambda}(0)\equiv 0\quad\text{ forall }\quad\ r\geq 4$ (1.5.14)
Proof. Write $22(r-1)=2r-p+p-$ 2, with $2(2r-p)=2r$, which implies $r=p$. Then, by (1.5.9) and the Schwarz inequality,
[…]
$\Box$
Remark 1.8.
In [CasJL76], the authors suggest in their Theorem 6.2 that infinite divisibility implies a Gaussian ch.f., under the sole assumption of the existence of its thermodynamic limit. Proposition  1.5 proves, however, the contrary, viz. that the ch.f. is Non-Gaussian for any finite volume, and, thus, Gaussian behavior may arise only in the thermodynamic limit. We show in the forthcoming Theorem 2.6 that this indeed happens, under certain assumptions.
The alternative proof sketched by the authors in a footnote, based on a result of G.G. Hegerfeldt ([Heg75]; the reference given in [CasJL76] is incorrect), would be indeed valid if $\tilde{\Phi}_{\Lambda,T}(t)$ were itself infinitely divisible, not only its reciprocal (1.5.7). This is an open problem.
2 The connected two-point function ($r=2$) in the critical case
2.1 The framework
We now turn to the framework introduced by one of us (M.R.) in [MR02], which is complementary to that introduced by Verbeure and collaborators, of which a comprehensive version is found in his book ([Ve11], see, in particular, Chap. 6, and the references given there). One different aspect in [MR02], which will be of central importance, is its use of a smoother version of fluctuation operators and Fourier transform methods. For a function $f$ defined on $\mathbb{Z}^{\nu}$, define its Fourier transform $\hat{f}$ as (see [BRo297], p.253):
$f(x)=(2\pi)^{-\nu}\int_{\begin{subarray}{c}-\pi<k_{i}\leq\pi;\\ i=1,\ldots,\nu\end{subarray}}d^{\nu}k\exp\left(ik\cdot x\right)\hat{f}(k)$ (2.1.1)
with
$\hat{f}(k)=\sum_{y\in\mathbb{Z}^{\nu}}\exp\left(-ik\cdot y\right)f(y)$ (2.1.2)
That is, the momenta $k$ are defined modulo the Brillouin zone $-\pi<k_{i}\leq\pi,i=1,\ldots,\nu$. Let
$f_{R}(x)\equiv f(\left|x\right|/R)$ (2.1.3)
with
$f(s)=\left\{\begin{array}[]{lll}1&\text{if}&0\leq s\leq 1\\ 0&\text{if}&s\geq 2\end{array}\right.$ (2.1.4)
be, for each $0<R<\infty$, a smooth (infinitely differentiable) function. We shall choose for $\Lambda$ a ball centered at the origin with radius $R$, let $R\rightarrow\infty$, and use (2.1.1), (2.1.2), with $f\rightarrow f_{R}$, given by (2.1.3) and (2.1.4) as the corresponding smooth approximation. It has the property (Lemma 2.6 of [MR02]):
$\hat{f_{R}}(k)=cR^{\nu}\hat{f}(Rk)$ (2.1.5)
where $c$ is a positive constant. We omit the dependence of $c$ on $R$, because it is uniformly bounded above and below by constants independent of $R$, the latter strictly positive (i.e. the r.h.s. of (2.1.5) is asymptotically correct as $R\rightarrow\infty$): according to the estimate, one of these two constants is employed. The smooth version of the normalized fluctuation operator (1.1.22) will be denoted by $F_{R}^{\alpha}(A)$, where
$F_{R}^{\alpha}(A)=R^{-\alpha}\sum_{x\in\mathbb{Z}^{\nu}}\left(A_{x}-\omega\left(A_{x}\right)\right)f_{R}(x)$ (2.1.6)
where the multiplicative constant $R^{-\alpha}$ corresponds to the asymptotic (for $R\rightarrow\infty$) order of the variance (1.1.23), with
$\alpha=\frac{\nu}{2}$ (2.1.7)
in the normal case, and
$\alpha>\frac{\nu}{2}$ (2.1.8)
in the anomalous case (1.1.21). In (2.1.6), $A$ stands for the observable, which will be $A=\sigma^{3}$ in the Ising case, and the reference to the thermal state $\omega_{T}$ is omitted, as in (1.1.22), not to burden the notation.
We now follow [MR02], Subsection 3.1. By translation invariance, the connected correlation functions $W_{\omega_{T},C}^{r}(x_{1},\ldots,x_{r})$ in (1.3.13), omitting now the reference to the temperature $T$ in $\omega$:
$W_{\omega,C}^{r}(x_{1},\cdots,x_{r})=W_{\omega,C}^{r}(x_{1}-x_{2},\cdots,x_{r-1}-x_{r})$ (2.1.9)
Introducing the variables
$y_{i}=x_{i}-x_{i+1}\ ,\qquad q_{i}=\sum_{j=1}^{i}k_{i}\ ,\quad i\leq r-1$ (2.1.10)
we now regard $W_{\omega,C}^{r}$ as function of the variables $\{y_{i}\}$, with $i\leq r-1$, and the Fourier transform $\widehat{W}_{\omega,C}^{r}$ as a function of the variables $q_{i};1\leq i\leq r-1$. We now pose
Assumption B $W_{\omega,C}^{r}$ is an $l^{1}$- function (i.e., absolutely summable) of the variables $y_{i}$, $i=1,\ldots,r$. As a consequence, $\widehat{W}_{\omega,C}^{r}$ is a continuous and uniformly bounded function of the variables $q_{i}$, $11\leq i\leq r-$ 1.
2.2 The case $0<T<T_{c}$
Theorem 3.3 of [MR02] may now be stated; we briefly sketch the proof for the reader’s convenience, because the argumements will be of major importance when dealing with the critical point:
Theorem 2.1.
Under Assumption B,
$0\leq\lim_{R\rightarrow\infty}\omega(F_{R}^{\nu/2}(A)^{r})<\infty\text{ \quad\ if \quad}r=2\ ,$ (2.2.1)
whereas
$\lim_{R\rightarrow\infty}\omega(F_{R}^{\nu/2}(A)^{r})=0\text{ \quad\ if \quad}r>2\ .$ (2.2.2)
Proof. By (2.1.5) and the definition (2.1.6),
$\omega(F_{R}^{\nu/2}(A)^{r})=\mbox{ const. }R^{r\nu/2}I$ (2.2.3)
where
$I=\int_{\begin{subarray}{c}|q_{i}|\leq\pi,\\ i=1,\ldots,r-1\end{subarray}}\prod_{i=1}^{r-1}dq_{i}\widehat{W}_{\omega,C}^{r}(q_{1},\dots,q_{r-1})J(\{Rq_{i}\})$ (2.2.4)
and
$J(\{Rq_{i}\})=\hat{f}(Rq_{1})\dots\hat{f}\left(-R(q_{1}+\cdots+q_{r-1})\right)$ (2.2.5)
By (2.2.3), (2.2.4) and (2.2.5), introducing the change of variable $q^{\prime}=qR$,
$\omega(F_{R}^{\nu/2}(A)^{r})=\mbox{ const. }R^{r\nu/2}R^{-(r-1)\nu}\int_{\begin{subarray}{c}|q_{i}^{\prime}|\leq R\pi;\\ i=1,\ldots,r-1\end{subarray}}J(\{q_{i}^{\prime}\})U(\{q_{i}^{\prime}\})$ (2.2.6)
where
$U(\{q_{i}^{{}^{\prime}}\})\equiv\widehat{W}_{\omega,C}^{r}(q_{1}^{\prime}/R,\ldots,q_{r-1}^{\prime}/R)$ (2.2.7)
By assumption (see (2.1.3), (2.1.4)), the $f$ functions are smooth and of compact support. Therefore $\hat{f}(k)$ is uniformly bounded in $k\in\mathbb{R}^{\nu}$ by a fixed function of fast decrease at infinity in $k$-space. As a consequence, Assumption B, (2.2.3) et seq. and the Lebesgue dominated convergence theorem yield (2.2.1) and (2.2.2).
$\Box$
What about the applicability of Theorem 2.1 to the Ising model? We have:
Theorem 2.2.
If $A=\sigma^{3}$ in (2.1.6), Assumption B holds for the Ising model with n.n. interactions (1.3.2). Consequently, a.) (2.2.1) holds if $\omega=\omega_{T}$ is any ergodic thermal state and $0\leq T<T_{c}$; b.) if $0\leq T\leq T_{0}<T_{c}$ and $T_{0}$ is sufficiently small, (2.2.2) holds as well. In (2.2.1),
$0<\lim_{R\rightarrow\infty}\omega_{T}(F_{R}^{\nu/2}(\sigma^{3})^{2})$ (2.2.8)
if and only if
$0<T<T_{c}$ (2.2.9)
Proof. The first assertion follows from [DCGR20], the second from ([BLOC98], Theorem 2, p. 431). By translation invariance of $\omega_{T}$ and Assumption B,
$\lim_{R\rightarrow\infty}\omega_{T}(F_{R}^{\nu/2}(\sigma^{3})^{2})=\sum_{x\in\mathbb{Z}^{\nu}}(\omega_{T}(\sigma_{x}^{3}\sigma_{0}^{3})-\omega_{T}(\sigma_{0}^{3})^{2})$
By Griffiths’ second inequality (1.3.17), each summand above is non-negative, and, therefore, restricting the sum to its value at $x=0$ we obtain
$\lim_{R\rightarrow\infty}\omega_{T}((F_{R}^{\nu/2}(\sigma^{3})^{2}))\geq 1/4-\omega_{T}(\sigma_{0}^{3})^{2}$
Again by Griffiths’ second inequality, the magnetization $\omega_{T}(\sigma_{0}^{3})$ is a decreasing function of the temperature, which tends to $1/2$ as $T\rightarrow 0$, leading to (2.2.8) and (2.2.9). The “only if” part follows from the fact that for $T=0$ each ergodic component is an (infinite) product state and therefore each term $\omega_{T}(\sigma_{x}^{3}\sigma_{0}^{3})-\omega_{T}(\sigma_{0}^{3})^{2}$ in the first sum above is identically zero.
$\Box$
Remark 2.1.
Case b.) was considerably extended by Newman ([Ne79], [Ne80]) to the whole region $0\leq T<T_{c}$, by a completely different and very elegant approach.
2.3 The case of anomalous fluctuations: the critical (ergodic) state for $r=2$ and $T=T_{c}$
We now come back to (2.1.6) and (2.1.9), assuming henceforth that $\omega=\omega_{T_{c}}$:
$\omega(F_{R}^{\alpha}(\sigma^{3})^{2})=R^{-2\alpha}\sum_{x,y\in\mathbb{Z}^{\nu}}f_{R}(\left|x\right|)f_{R}(\left|y\right|)W_{\omega,C}^{2}(\left|x-y\right|)$ (2.3.1)
where, by translation invariance of $\omega$,
$W_{\omega,C}^{2}(\left|x\right|)=\omega((\sigma_{x}^{3}-\omega(\sigma_{x}^{3}))(\sigma_{0}^{3}-\omega(\sigma_{0}^{3})))$ (2.3.2)
The function $W_{\omega,C}^{2}$ is a uniformly bounded function on $\mathbb{Z}^{\nu}$. Let $\mathbb{T}_{\nu}=(-\pi,\pi]^{\nu}$ be the $\nu$- dimensional torus and $\mathcal{D}(\mathbb{T}_{\nu})$ denote the space of all functions $\tilde{\phi}$ on $\mathbb{T}_{\nu}$ such that
$\tilde{\phi}\in C^{\infty}(\mathbb{R}^{\nu})\qquad\text{ with }\mathit{\qquad}\tilde{\phi}(k_{1},\ldots,k_{n})=\phi(\exp(ik_{1}),\ldots,\exp(ik_{n}))$ (2.3.3)
(see [Ru73], p.190, exercise 22). If $\tilde{\phi}\in\mathcal{D}(\mathbb{T}_{\nu})$, then $\sum_{x\in\mathbb{Z}^{\nu}}(1+\left|x\right|^{N})\left|\phi(x)\right|^{2}<\infty$ for all integers $N<\infty$. These norms define a Fréchet space topology on $\mathcal{D}(\mathbb{T}_{\nu})$ and any function such as $W_{\omega,C}^{2}$ satisfying
$\left|W_{\omega,C}^{2}(\left|x\right|)\right|\leq\text{ const. }(1+\left|x\right|)^{M}$ (2.3.4)
(there is an obvious misprint in [Ru73]), for some positive integer $M$, then it is defined as a distribution on $\mathbb{Z}^{\nu}$ by the equation:
$W_{\omega,C}^{2}(\phi)\equiv\sum_{x\in\mathbb{Z}^{\nu}}W_{\omega,C}^{2}(x)\phi(x)$ (2.3.5)
This distribution is of positive type, i.e.,
$W_{\omega,C}^{2}(\bar{\phi}_{-}\ast\phi)\geq 0\ ,\qquad\forall\phi\in\mathbb{Z}^{\nu}$ (2.3.6)
where
$(f\ast g)(x)\equiv\sum_{y\in\mathbb{Z}^{\nu}}f(x-y)g(y)$ (2.3.7)
and
$\phi_{-}(x)\equiv\phi(-x)$ (2.3.8)
Notice that $W_{\omega,C}^{2}$ satisfies (2.3.4), being uniformly bounded. By a simplified version of the Bochner-Schwartz theorem (Theorem IX.10 of [RS75]), and a simplified version of the argument at the bottom of p.14 of [RS75], we find
Theorem 2.3.
$W_{\omega,C}^{2}$ is the Fourier transform of a positive, finite measure $\mu$ on the torus $\mathbb{T}_{\nu}$.
According to the above theorem, we may thus write
[…]
where $\mu$ is a positive, finite measure on $\mathbb{T}_{\nu}$, which has the decomposition
$\mu=\mu_{\mathrm{a.c.}}+\mu_{\mathrm{p.p.}}+\mu_{\mathrm{s.c.}}$ (2.3.10)
([RS72], Theorems I.13 and I.14), where $\mu_{\mathrm{p.p.}}$ is pure point, $\mu_{\mathrm{a.c.}}$ absolutely continuous and $\mu_{\mathrm{s.c.}}$ singular continuous with respect to Lebesgue measure on the torus. By Theorem 2.3, $\mu_{\mathrm{a.c.}}$, $\mu_{\mathrm{p.p.}}$ and $\mu_{\mathrm{s.c.}}$ are all positive, finite measures on the torus.
In the proof of Theorem 2.1, we have used that $\widehat{W}_{\omega,C}^{2}$ is a continuous and uniformly bounded function on the torus. Comparing with (2.3.9), this means that
$d\mu(p)=\widehat{W}_{\omega,C}^{2}(p)d^{\nu}p$ (2.3.11)
where
$\widehat{W}_{\omega,C}^{2}(p)\text{ is a uniformly bounded continuous function on }\mathbb{T}_{\nu}$ (2.3.12)
We now assume that this condition does not hold, which implies that $W_{\omega,C}^{2}$ is not an $l^{1}$-function, in agreement with Criterion A. We now have:
Theorem 2.4.
$\widehat{W}_{\omega,C}^{2}$ is a positive definite distribution on the torus $\mathbb{T}_{\nu}$. Write
$d\mu_{\mathrm{a.c.}}(p)=g(p)d^{\nu}p$ (2.3.13)
but, in contrast to (2.3.12), assume that
$g\text{ is not a uniformly bounded continuous function on }\mathbb{T}_{\nu}$ (2.3.14)
Then the observable $\sigma^{3}$ exhibits anomalous fluctuations in the state $\omega$, with, in (1.1.21),
$\alpha=\nu/2+\rho/2$ (2.3.15)
where
$0<\rho<\nu$ (2.3.16)
If the above equations (2.3.15) and (2.3.16) hold, the contributions of $\mu_{\mathrm{p.p.}}$ and $\mu_{\mathrm{s.c.}}$ vanish as $R\rightarrow\infty$.
Proof. The first assertion follows from (2.3.2), the definition (2.3.6) and the positivity of the state, i.e.,
$\omega(A^{\ast}A)\geq 0\ ,\qquad\forall A\in\mathcal{A}$ (2.3.17)
We first assume that $\mu_{\mathrm{p.p.}}=0$ and $\mu_{\mathrm{s.c.}}=0$.
We may write the contribution of the absolutely continuous part (2.3.11) to the r.h.s. of (2.3.1), by the convolution theorem as
$I_{R}$$\equiv$ $R^{-2\alpha}R^{\nu}\int_{\left|p\right|\leq\pi}\hat{f}(Rk)\hat{f}(-Rk)g(k)d^{\nu}k$ (2.3.18)
$=$ $R^{-2\alpha}\int_{\left|k\right|\leq\pi R}\left|\hat{f}(k^{\prime})\right|^{2}g(k^{\prime}/R)d^{\nu}k^{\prime}$
If $g$ is not a uniformly bounded continuous function, it must exhibit at some point of the torus (without loss, at the origin) an integrable singularity, i.e.,
$g(k^{\prime}/R)=O(\left(\left|k^{\prime}\right|/R\right)^{-\rho})\qquad\text{as}\qquad\left|k^{\prime}\right|\rightarrow 0$ (2.3.19)
with (2.3.16) valid, because the measure $\mu_{\mathrm{a.c.}}$ is finite by Theorem (2.3). By (2.1.3) and (2.1.4), $\hat{f}$ is smooth and rapidly decreasing, and thus, by (2.3.16), the integral
$\int_{\mathbb{R}^{\nu}}\left|\hat{f}(k^{\prime})\right|^{2}\left|k^{\prime}\right|^{-\rho}d^{\nu}k^{\prime}<\infty$ (2.3.20)
Let $\alpha$ be defined by (2.3.15). Then, by (2.3.18), (2.3.19) and (2.3.20), $\lim_{R\rightarrow\infty}I_{R}$ exists and, by the first inequality in (2.3.16), the observable $A=\sigma^{3}$ has anomalous fluctuations in the state $\omega$, according to the definition (1.1.21).
In order to show the last assertion, observe that the contributions of $\mu_{\mathrm{p.p.}}$ and $\mu_{\mathrm{s.c.}}$ to the r.h.s. of (2.3.1) equal
$\lim_{R\rightarrow\infty}R^{-2\alpha}R^{\nu}\int_{|k|\leq\pi}\left|\hat{f}(Rk)\right|^{2}\left(d\mu_{\mathrm{p.p.}}(k)+d\mu_{\mathrm{s.c.}}(k)\right)=0$
by (2.2.8), (2.2.9), and the bound
$\left|\hat{f}(Rk)\right|\leq\mbox{ const. }$.
with the constant above independent of $R$, which results from (2.1.5).
$\Box$
Let $f\in C_{0}^{\infty}(\mathbb{R}^{\nu})$. Then ([LL00], Theorem 5.9: Fourier transform of $\left|x\right|^{\rho-\nu}$), with $c_{\rho}\equiv\pi^{-\rho/2}\Gamma(\rho/2)$,
[…]
where the superscript denotes the inverse of the Fourier transform. We have, however, replaced their $\alpha$ by our $\rho$, which satisfies (2.3.16), i.e., their assumption. Our conventions for the Fourier transform differ from those in [LL00], so that the above formulas may differ from those in our notation by constants which do not alter the final statement. If we now set
$\rho=2-\eta$ (2.3.22)
we obtain that
$W_{\omega,C}^{2}(|x|)=\left|x\right|^{-(\nu-2+\eta)}$ (2.3.23)
satisfies (2.3.21). For $\rho>\nu/2$, the r.h.s. of (2.3.21) is not in $L^{p}(\mathbb{R}^{\nu})$ for any $p\leq 2$, as the authors remark p. 131 of [LL00], which makes inclusion of an $f\in C_{0}^{\infty}(\mathbb{R}^{\nu})$ in (2.3.21) necessary. In the lattice case these considerations become simplified, but (2.3.19) shows that only the asymptotic behavior at large distances is an issue, and, accordingly, we replace (2.3.23) by the
Standard Assumption ([GML75], [CasJL78])
$W_{\omega,C}^{2}(|x|)=O(\left|x\right|^{-(\nu-2+\eta)})\qquad\text{ as \qquad}\left|x\right|\rightarrow\infty$ (2.3.24)
Remark 2.2.
The standard assumption is proved true if $\nu=2$ by a closed-form calculation [MW73], yielding
$\eta=0.25$ (2.3.25)
The following theorem is therefore complementary to Theorem 2.4:
Theorem 2.5.
Under the Standard Assumption (2.3.24), $W_{\omega,C}^{2}(|x|)$ is not an $l^{1}$- function, and the observable $A=\sigma^{3}$ has anomalous fluctuations in the state $\omega$, according to definition (1.1.21).
Proof. By (2.3.24) and Theorem 17.8 by Glimm and Jaffe [GJ81],
$\eta\leq 1$ (2.3.26)
and hence, by (2.3.24), $W_{\omega,C}^{2}(|x|)$ is not an $l^{1}$- function. By (2.3.24) and Griffiths’ second inequality (1.3.17), there exists a $R_{0}<\infty$ and two constants $c>0$ and $d<\infty$ such that
$c~{}\left|x\right|^{-(\nu-2+\eta)}\leq W_{\omega,C}^{2}(|x|)\leq d~{}\left|x\right|^{-(\nu-2+\eta)}\qquad\mbox{ for all }\qquad\left|x\right|\geq R_{0}$ (2.3.27)
We now write
$\alpha=(\nu+\rho)/2$ (2.3.28)
By (2.3.1) and (2.3.28),
$\omega(F_{R}^{\alpha}(\sigma^{3})^{2})=R^{-\nu}R^{-\rho}\sum_{x,y\in\mathbb{Z}^{\nu}}f_{R}(\left|x\right|)f_{R}(\left|y\right|)W_{\omega,C}^{2}(\left|x-y\right|)$ (2.3.29)
By (2.1.3), (2.1.4), (2.3.26), (2.3.27), (2.3.29) and Griffiths’ second inequality (1.3.17), we obtain
[…]
$=$ $cR^{-\rho}[(2R)^{2-\eta}-R_{0}^{2-\eta}]+o\left(R\right)$
By (2.3.26), choosing
$\rho=2-\eta$ (2.3.31)
the $R_{0}$- term in (2.3.30) does not contribute in the limit $R\rightarrow\infty$, and we obtain from (2.3.30)
$\lim_{R\rightarrow\infty}\omega((F_{R}^{\alpha}(\sigma^{3})^{2}))\geq c~{}2^{2-\eta}$ (2.3.32)
The second inequality in (2.3.27) yields similarly, with $\alpha$ and $\rho$ as in (2.3.28) and (2.3.31),
$\lim_{R\rightarrow\infty}\omega((F_{R}^{\alpha}(\sigma^{3})^{2}))\leq d~{}2^{2-\eta}$ (2.3.33)
By (2.3.32), (2.3.33) ,(2.3.28) and (2.3.31), the second assertion of the theorem follows.
$\Box$
Remark 2.3.
Equation (2.3.26) was proved in [GJ81], Theorem 17.8, under the sole assumption (2.3.24). It uses in an essential way the Lebowitz inequality (1.3.18), as well as the Simon-Lieb correlation inequality ([Si80], [Li80]).
Remark 2.4.
Theorems 2.4 and 2.5 together make Criterion A very plausible. Assumption (2.3.14) implies that $W_{\omega,C}^{2}(|x|)$ is not an $l^{1}$- function, and has, at the same time, an integrable singularity in its Fourier transform, which is related, in a certain precise sense previously explained, to the asymptotic form (2.3.24). On the other hand, if (2.3.24) is assumed, then $W_{\omega,C}^{2}(|x|)$ is not an $l^{1}$- function, and the observable $A=\sigma^{3}$ has anomalous fluctuations in the state $\omega$, according to Definition (1.1.21).
The above remark refers to the first inequality in (2.3.16) (given (2.3.15)). It is also important to notice that the second inequality in (2.3.16) also has a physical-mathematical content which is natural in the present approach. Indeed, setting (2.3.22) into (2.3.16), we find the requirement
$2-\eta<\nu$
or
$\nu-(2-\eta)>0$ (2.3.34)
Comparing with (2.3.24), we see that (2.3.34) is just the requirement of ergodicity (1.2.8) of the state $\omega$, which is a general result for quantum spin systems by [BRo186], Example 4.3.24, p.401.
2.4 The connected $r$-point functions in the critical case ($r\geq 4$ ) and the quasi-free nature of $\omega_{T}^{F}$
We now come back to (2.2.6) and (2.2.7), and pose, for all $r$, $A=\sigma^{3}$,
$\omega_{T_{c}}(F_{R}^{\nu/2+\rho/2}(\sigma^{3})^{r})=\text{const. }R^{r\nu/2}R^{-(r-1)\nu}R^{\rho^{\prime}(r-2)}$ (2.4.1)
where the “constant” is to be understood as in (2.1.5) et seq.. In (2.4.1) we introduce a new exponent $\rho^{\prime}$, for $r\geq 4$. This Ansatz has the general form of the scaling hypothesis VI, p.434, in [GML75]:
Scaling Assumption C
$\widehat{W}_{\omega_{T_{c}},C}^{r}(q_{1}^{\prime}/R,\ldots,q_{r-1}^{\prime}/R)=R^{\rho^{\prime}(r-2)}\widehat{W}_{\omega_{T_{c}},C}^{r}(q_{1}^{\prime},\ldots,q_{r-1}^{\prime})\qquad\text{ for \qquad}r\geq 4$ (2.4.2)
with an exponent $\rho^{\prime}$ satisfying
$0<\rho^{\prime}<\infty$ (2.4.3)
$\widehat{W}_{\omega_{T_{c}},C}^{r}$ is, in fact, a (Schwartz) distribution defined by
$\langle\widehat{W}_{\omega_{T_{c}},C}^{r},\hat{f}\rangle=\langle W_{\omega_{T_{c}},C}^{r},f\rangle\equiv\sum_{x_{1},\ldots x_{r}\in\mathbb{Z}^{\nu}}W_{\omega_{T_{c}},C}^{r}(x_{1}\ldots x_{r})f(x_{1},\ldots,x_{r})$
for functions $f$ of bounded support. Equation (2.4.2) corresponds to defining the scaled distribution $\widehat{W}_{\omega_{T_{c}},C}^{r,R}$ associated to $\widehat{W}_{\omega_{T_{c}},C}^{r}$ by
$\langle\widehat{W}_{\omega_{T_{c}},C}^{r,R},\hat{f}\rangle\equiv\langle\widehat{W}_{\omega_{T_{c}},C}^{r},\hat{f}_{R}\rangle$
where
$\hat{f}_{R}(q_{1}^{\prime},\ldots q_{r-1}^{\prime})\equiv R^{\nu(r-1)}f(Rq_{1}^{\prime},\ldots,Rq_{r-1}^{\prime})$
with $|q_{i}^{\prime}|\leq\pi,i=1,\cdots,r-1$. Finally, equation (2.4.2) should be written precisely as
$\langle\widehat{W}_{\omega_{T_{c}},C}^{r},\hat{f}_{R}\rangle=R^{\rho^{\prime}(r-2)}\langle\widehat{W}_{\omega_{T_{c}},C}^{r},\hat{f}\rangle$
Note also that the condition $\rho^{\prime}>0$ corresponds, for $r\geq 4$, to the analogue of the first of (2.3.16) for $r=2$. For $r=2$ in (2.4.1), the r.h.s. of (2.4.1) is $O(1)$ by Theorems 2.4 and 2.5.
Theorem 2.6.
Assume that
[…]
and, moreover, the limit agrees with its LLRO version, that is to say, one may replace in (2.4.4), $\omega_{T_{c}}$ by $\omega_{R,T_{c}}$. Then, under the Scaling Assumption C,
$\lim_{R\rightarrow\infty}\omega_{T_{c}}((F_{R}^{\nu/2+\rho/2}(\sigma^{3})^{r})=0\quad\text{ for }\quad r\geq 4$ (2.4.5)
and thus the fluctuation state $\omega_{T_{c}}^{F}$ is quasi-free.
Proof. By (2.4.1) and Scaling Assumption C,
$\omega_{T_{c}}(F_{R}^{\nu/2+\rho/2}(\sigma^{3})^{r})=\text{const. }R^{r(\rho^{\prime}-\nu/2-\rho/2)}R^{\nu-2\rho^{\prime}}$ (2.4.6)
Assume, first, that
$20<\rho^{\prime}<\alpha=(\nu+\rho)/$ 2 (2.4.7)
Then, by (2.4.6) and (2.4.7), for $r$ sufficiently large,
$\lim_{R\rightarrow\infty}\omega_{T_{c}}(F_{R}^{\nu/2+\rho/2}(\sigma^{3})^{r})=0$ (2.4.8)
and, hence, by the assumption of agreement with the LLRO version,
$\lim_{R\rightarrow\infty}\omega_{R,T_{c}}(F_{R}^{\nu/2+\rho/2}(\sigma^{3})^{r})=0\mbox{ for sufficiently large }r$ (2.4.9)
It follows from (2.4.9) and Proposition  (1.6) that
$\lim_{R\rightarrow\infty}\omega_{R,T_{c}}(F_{R}^{\nu/2+\rho/2}(\sigma^{3})^{r})=0\mbox{ for all }r\geq 4$ (2.4.10)
and therefore the fluctuation state is quasi-free. If, on the other hand,
$\rho^{\prime}=\alpha$ (2.4.11)
then $\nu-2\rho^{\prime}=2(\nu/2-\rho^{\prime})=-\rho<0$ and (2.4.5) holds as well. The last case, $\rho>\alpha$ implies that the r.h.s of (2.4.6) diverges, as $R\rightarrow\infty$, for sufficiently large $r$, contradicting the assumption (2.4.4).
$\Box$
Remark 2.5.
It is the necessity of using a definite normalization for all $r$ in (2.4.4) – the only one which guarantees the existence of the thermodynamic limit for $r=2$, due to the existence of anomalous fluctuations – which leads to a quasi-free fluctuation state in the above theorem; in particular, the fact that the case of equality (2.4.11) satifies (2.4.5).
Remark 2.6.
By Theorem 2.6 and Proposition 1.6 the fluctuation state undergoes a basic structural change, from interacting to quasi-free, in the passage to the thermodynamic limit. This is different from previous references to structural changes of states, which concerned states of infinite systems: one exception is the change from pure to mixed state in the quantum theory of measurement in [Wre123], which also occurred in the passage to the thermodynamic limit. Both transformations, therefore, occur naturally, and should be accounted for in a general theory, see Section 4.
3 Long-range interactions: universality and its violations
It is well-known that the critical exponents or indices depend only on rather general qualitative features of physical systems, such as dimensionality, symmetry and range of interaction: this is the universality of critical phenomena [Fis67]. We shall be concerned only with the Ising universality class, which concerns only a discrete symmetry (the spin-reflection symmetry), but consider in this section long-range interactions. Indeed, the short vs. long range character of the interaction may be of decisive importance in determining whether a phase transition of a given type (e.g., first-order) takes place in a system, e.g. for Ising systems in dimension $\nu=1$ no first-order phase transition is possible if the interaction is short-range; in order to have one, we must consider long-range interactions, to which we now turn.
3.1 Brief review of results and problems
Concerning long-range interactions, we shall limit ourselves to a prototype, the Dyson [Dy69] model - referred to as $D_{\alpha}$-model. It is given by (1.2.1) with $\delta=0$ and
$j(\left|x\right|)=\frac{1}{\left|x\right|^{\alpha}}$ (3.1.1)
in dimension
$\nu=1$ (3.1.2)
For
$21<\alpha<$ 2 (3.1.3)
it was proved by Dyson that the model exhibits a ferromagnetic phase transition (for an independent proof, see [FILS78]). It has been proved (see [AF88], Prop. 2.1) that several critical exponents, including $\beta$ and $\gamma$ in (1.1.17) and (1.1.18), assume their mean-field values, viz.
$\beta=\frac{1}{2}$ (3.1.4)
and
$\gamma=1$ (3.1.5)
(there are others, see [AF88]), as soon as
$1<\alpha\leq\frac{3}{2}$ (3.1.6)
This is not quite (3.1.3), but it is believed that the result is correct, in agreement with the predictions of Bleher and Sinai ([BleSi73], [BleSi75], see also [Sew86], Sec. 5.4.3 for a pedagogic outline) for the Dyson hierarchical model. It is of interest that Lohmann, Slade and Wallace [LSW17] have proved that the exponent $\eta$ keeps its mean-field value even for $\alpha>3/2$. We come back in Section 4 to a more detailed analysis of the predictions of Bleher and Sinai.
This universality – the independence of the critical indices on the rate $\alpha$ of fall-off of the potential (3.1.1) – need not happen for all “critical indices”, in particular not for those related to the dynamics. This issue – the breaking of universality for “dynamical critical indices” – has been seldom studied: it is the subject of the next subsection.
Remark 3.1.
Regarding the interval (3.1.6) in which critical exponents $\beta$ and $\gamma$ assume their mean-field values (3.1.4) and (3.1.5), it seems, at first sight, that no dimension may be attached to the power decay $\alpha$ of the interaction in $D_{\alpha}$-model. However, since the hierarchical model has been introduced by Dyson to establish long range order in the $D_{\alpha}$-model; since the spectral dimension $d$ of the hierarchical Laplacean $\Delta^{\mathrm{hier.}}$ (see e.g. [BAHM02]) is able to distinguish whether a random walk generated by $\Delta^{\mathrm{hier.}}$ is recurrent if $d\leq 2$ or transient if $d>2$; the function $\alpha(d)=(d+2)/d$ deduced from these facts establishes that $\alpha(\infty)=1$, $\alpha(2)=2$ and $\alpha(4)=3/2$. So, $21<\alpha\left(d\right)<$ 2 as $d\in\left(2,\infty\right)$ and the interval (3.1.6) refers to $d\geq 4$, the range for which the fluctuation in the Ising model at critical point are Gaussian and the critical exponents are classical. See also [ADCV15] Proposition 1.3, Corollaries 1.4 and 1.5 and references given there for a closely related viewpoint using infrared bound and substantial argument based on the continuity of order parameter.
3.2 Aspects of dynamics: violations of universality in problems of energetic and dynamic stability
Up to now we have made little reference to the subject of dynamics. The dynamics of fluctuations is of more importance as in the work of Wehrl [We71] which proved the important fact that the “true” fluctuations (i.e., corresponding to zero momentum) do not satisfy the KMS condition, suggesting the quasi-free nature of the fluctuation states for all temperatures (see Remark 1.7). How do the other macroscopic observables – the intensive observables – behave in this connection, i.e., in a non-equilibrium framework?
In this section we consider the dynamics of intensive observables, viz. the mean magnetization. It is shown that this dynamics is able to distinguish the two universality classes – short range and long-range – through the different rates of decay of unstable states in the system, viz. macroscopic “liquid drops”. Even more interestingly, these effects are able to distinguish the several types of (nonuniversal) behavior within one class (e.g., a class of long-range interactions), a topic which has been very seldom touched upon.
A heuristic, but illuminating way to understand the difference of behavior of short and long-range interactions was proposed by Haag in his seminal paper on the mathematical structure of the BCS model [Ha62]. Consider, for simplicity, the ground state of a ferromagnet (e.g., Ising or Heisenberg) with spins oriented along the $+3$-direction (say). He argued that the energy necessary to build a “liquid droplet” of oppositely oriented spins (along the $-3$-direction) of large volume $V$ is of larger order in $V$ for long-range interactions than for those of short range. Thus, long-range interactions stabilize the system: we shall call this “freezing” property of long-range interactions “Haag’s principle”. When applying Haag’s principle to particular models, it will be possible to make more precise the conditions of its validity.
A prototype of models of long-range interactions is Dyson’s $D_{\alpha}$ model. The energy $E$ necessary to create a negatively polarized droplet of length $N$ in a positively polarized state equals (see also [Sew86], p.126, which we follow here):
$E=\text{const.}N^{2-\alpha}$ (3.2.1)
In the critical region (3.1.3), this energy is of an order of magnitude greater than that which is necessary for short range interactions (e.g., bounded by a constant in the case of finite range). This is an expression of Haag’s principle for the Dyson model, whose validity is, however, restricted to the values of $\alpha$ corresponding to the phase-transition region. This is not an accident: (3.2.1) is the central ingredient in an energy-entropy estimate which may even be used to prove the existence of a phase transition in the model (see [Sew86], p. 126 and references given there). A proof of phase transitions in the Dyson model along these lines was given in [CFMP05].
If one considers now a state consisting of a large droplet of spins polarized along the $1$-direction, immersed in a background of spins polarized along the $3$-direction, the energy fluctuation $\Delta E$ of the Dyson Hamiltonian in this state is of the order of the r.h.s. of (3.2.1). This state will be unstable and decay with a half-life $\Gamma$ given by the time-energy uncertainty principle $\Delta E\cdot\Gamma\approx\hbar$ (in the case of a Markovian approach to equilibrium , i.e., of type $\exp(-t/\Gamma)$), otherwise, instead of $\Gamma$, one would have a quantity called “sojourn time” in the state (see, e.g., [MWB13], Theorem 3.17, p. 81, and references given there). From this heuristics and (3.2.1), we are led to conclude: the rate of decay of such a state is the greater, the longer the range of interaction, i.e, the closer the parameter $\alpha$ is to one, inside the open interval $(1,2)$. We shall call this heuristic principle, for brevity, “Haag’s principle”, because it is a direct consequence of the rise of energetic stability of large droplets when the range of interaction increases. If, in (1.2.1), we consider the Ising limit, i.e., set $\delta\equiv 0$, we obtain the generalized Ising model (gIm) studied by Emch [Em66] and Radin [Ra70] (we changed the overall sign to agree with Radin’s notation)
$H_{\Lambda}=\frac{1}{2}\sum_{x,y\in\Lambda}j(|x-y|)\sigma_{x}^{3}\sigma_{y}^{3}$ (3.2.2)
For a class of initial states, characterized in [Ra70], of which the product states
$\omega_{p,q}\equiv\prod_{x\in\mathbb{Z}^{\nu}}\left(p~{}\left|\pm\right)_{x}+q~{}\left|\pm\right)_{x}^{\prime}\right)$ (3.2.3)
are prototypes, where $p$ and $q$ are (without loss) non-negative and
$\sigma_{x}^{1}~{}\left|\pm\right)_{x}=\pm~{}\left|\pm\right)_{x}\qquad\text{ and\qquad\ }\sigma_{x}^{2}~{}\left|\pm\right)_{x}^{\prime}=\pm~{}\left|\pm\right)_{x}^{\prime}\qquad\text{ and \qquad}p^{2}+q^{2}=1$ (3.2.4)
the approach to equilibrium is explicit [Em66]:
$\omega_{p,q}(\tau_{t}(\sigma_{x}^{1}))=\pm p\prod_{y\in\mathbb{Z}^{\nu}}\cos^{2}(2J(\left|y\right|)t)\qquad\text{ for any }x\in\mathbb{Z}^{\nu}$ (3.2.5)
with a corresponding relation for the $2$-component. By the stability condition (1.2.4),
$\lim_{\left|x\right|\rightarrow\infty}j(\left|x\right|)=0\,,$
and, in addition, the above infinite product converges absolutely. Let, now, $\nu=1$, and define a subclass of models of the gIm:
For Dyson’s $D_{\alpha}$ model, the infinite product in (3.2.5), relevant to the decay of the magnetization along the $1$- or $2$-axis, is given by
$\mathrm{Cl}_{1;\alpha}(t)\equiv\prod_{j\in\mathbb{N}}\cos(j^{-\alpha}t)$ (3.2.6)
where $\mathrm{Cl}$ denotes the Cloitre function introduced in [AlKi17]. We have the following lemma due to Radin ([Ra70], Lemma4, p. 2954), whose proof we reproduce for the reader’s convenience, because it “explains” part of the forthcoming theorem of Albert and Kiessling (Theorem 5.1 of [AlKi17]):
Lemma 3.1.
If $\alpha>1$, $\exists~{}c>0$ such that
$\left|\mathrm{Cl}_{1;\alpha}(t)\right|\leq\exp(-c|t|^{1/\alpha})$ (3.2.7)
Proof. For $10<x<$ 1, we have $0<\cos x<1-cx^{2}$, for some $c>0$, and $1-x\leq\exp(-x)$. Therefore, for $t>0$, we have
$|\mathrm{Cl}_{1;\alpha}(t)|$$\leq$ $\prod\nolimits_{j>t^{1/\alpha}}\left|\cos(t/j^{\alpha})\right|$
$\leq$ $\prod\nolimits_{j>t^{1/\alpha}}\left(1-ct^{2}/j^{2\alpha}\right)$
$\leq$ $\prod\nolimits_{j>t^{1/\alpha}}\exp\left(-ct^{2}/j^{2\alpha}\right)=\exp\left(-c\sum\nolimits_{j>t^{1/\alpha}}t^{2}/j^{2\alpha}\right)$
By integral approximation, $\sum_{j>t^{1/\alpha}}t^{2}/j^{2\alpha}\geq t^{1/\alpha}$, and transition to negative $t$ finally yields the full result.
$\Box$
The following special case of Theorem 5.1 of [AlKi17] considerably improves Lemma 3.1:
Theorem 3.2.
Let
$\alpha>\frac{1}{2}$ (3.2.8)
Then, $\forall t\in\mathbb{R}$,
$\mathrm{Cl}_{1;\alpha}(t)=\exp(-C_{1;\alpha}|t|^{1/\alpha})F_{1;\alpha}(|t|)$ (3.2.9)
where
$\left|F_{1;\alpha}(|t|)\right|\leq\exp(K_{1;\alpha}\left|t\right|^{1/(\alpha+1)})\ ,\qquad\text{ for some constant }K_{1;\alpha}>0$ (3.2.10)
and where
$C_{1;\alpha}=-\frac{1}{\alpha}\int_{0}^{\infty}\log\left|\cos\xi\right|\frac{1}{\xi^{1+1/\alpha}}d\xi$ (3.2.11)
The proof of the above theorem in [AlKi17] follows in part Theorem 1 of [Ki13]. Since increasing the range of interaction in the Dyson model in the range (3.1.3) means varying $\alpha$ from $2$ to $1$, in the open interval, we have the following
Corollary 3.3.
$\left|\omega_{p,q}(\tau_{t}(\sigma_{x}^{1}))\right|=O(\exp(-c\left|t\right|^{1/\alpha})\ ,\qquad\text{for some }c>0\text{ as }\left|t\right|\rightarrow\infty\ .$ (3.2.12)
(3.2.12) defines a “dynamical critical exponent” $1/\alpha$, which satisfies Haag’s principle and violates universality.
The question may now be posed whether the mean-field limit describes the decay of the states $\omega_{p,q}$ for large range of interaction.
We now show that there exists a (trivial) exact solution of the non-linear system derived in Appendix A of [LvH78], of which a special case is the mean-field limit of the model of the previous section, with the initial state (3.2.3), (3.2.4) and that, as a consequence, the mean-field limit does not describe the decay of the states $\omega_{p,q}$ for large range of interaction.
We consider the (mean-field) Hamiltonian
$H_{\Lambda}^{\mathrm{m.f.}}\equiv\frac{1}{|\Lambda|}\sum_{x,y\in\Lambda}H^{\mathrm{m.f.}}(x,y)$ (3.2.13)
where
$H^{\mathrm{m.f.}}(x,y)=a\sigma_{x}^{1}\sigma_{y}^{1}+b\sigma_{x}^{2}\sigma_{y}^{2}+c\sigma_{x}^{3}\sigma_{y}^{3}$ (3.2.14)
The above corresponds to the Hamiltonian (3.2.1) for a van der Waals spin system in reference [LvH78]. We define, as in (A.1) of ([LvH78], Appendix A), the time-dependent spin operator (changing to the notation $S(x;t)\rightarrow\sigma(x;t)$)
$\sigma^{i}(x;t)\equiv\exp(iH_{\Lambda}t)\sigma^{i}(x;0)\exp(-iH_{\Lambda}t)\qquad\text{ for }i=1,2,3$ (3.2.15)
and, as in [LvH78], the time-dependent space-average of $\sigma^{i}(x;t),i=1,2,3$ by
$\eta_{\Lambda}(\sigma^{i};t)\equiv\frac{1}{|\Lambda|}\sum_{x\in\Lambda}\sigma^{i}(x;t)$ (3.2.16)
We have the following immediate consequence of the methods of [LvH78], Appendix A:
Lemma 3.4.
In the representation defined by the state $\omega_{p,q}$, defined by (3.2.3) and (3.2.4), the following strong limit exists:
$\lim_{\Lambda\nearrow\infty}\eta_{\Lambda}(\sigma^{i};t)=M^{i}(t)\mathbf{1}\qquad\text{ for }i=1,2,3$ (3.2.17)
where the $M^{i}(t),i=1,2,3$, satisfy the system of nonlinear differential equations on $\mathbb{R}^{3}$:
$\dot{M}^{1}(t)=2(b-c)M^{2}(t)M^{3}(t)$ (3.2.18)
$\dot{M}^{2}(t)=2(c-a)M^{1}(t)M^{3}(t)$ (3.2.19)
$\dot{M}^{3}(t)=2(a-b)M^{1}(t)M^{2}(t)$ (3.2.20)
Moreover, equation (3.2.17) is the precise analog of the l.h.s. of equation (3.2.5) when, in $H_{\Lambda}^{\mathrm{m.f.}}$ of (3.2.13) and (3.2.14), the special case $a=b=0$ is considered.
Proof. Concerning the mean-field model, $\omega_{p,q}$, as defined by (3.2.3), (3.2.4), being a product state, i.e., satisfying $\omega_{p,q}(AB)=\omega_{p,q}(A)\omega_{p,q}(B)$, $\forall A,B\in\mathcal{A}$, trivially satisfies the condition of ergodicity (1.2.8), and therefore, by Lemma 2, p. 402 of [LvH78], the strong limits in (3.2.17) exist. The rest follows by the Lemma p. 435 in [LvH78].
$\Box$
The following theorem shows that mean-field theory dramatically fails to describe the decay in Corollary 3.3:
Theorem 3.5.
Let, in (3.2.13), (3.2.14),
$a=b$ (3.2.21)
as well as
$M^{3}(0)=0$ (3.2.22)
Then, the system of nonlinear equations (3.2.18), (3.2.19), (3.2.20) on $\mathbb{R}^{3}$ has the unique, global solution given by
$M^{1}(t)\equiv\pm p\qquad\text{ together with \qquad}M^{2}(t)\equiv\pm q\text{\quad and \quad}M^{3}(t)\equiv 0$ (3.2.23)
that is, the mean magnetization remains “frozen” at its initial value (±$(\pm p,\pm q,0)$. The solution is Lyapunov stable.
Proof. By (3.2.21), (3.2.22) and (3.2.20), $M^{3}(t)\equiv 0$, $\forall t\in\mathbb{R}$. Inserting this in (3.2.18) and (3.2.19), we obtain (3.2.23).
The quantity $\Lambda(t)\equiv M^{1}(t)^{2}+M^{2}(t)^{2}+M^{3}(t)^{2}$ is a constant of the motion (Casimir element), even without the assumption (3.2.21), which guarantees that the system of equations (3.2.18),(3.2.19),(3.2.20) has a unique global solution, see [LvH78], Appendix A, and references given there.
The quantity $\Lambda$ is also a Lyapunov function for the system of nonlinear differential equations (3.2.18), (3.2.19), (3.2.20), satisfying $\dot{\Lambda}(t)\equiv 0$, and therefore the solution is Lyapunov-stable.
$\Box$
The above theorem implies, together with Lemma 3.4 and Corollary 3.3, that the mean-field version of (3.2.2) violates Haag’s principle in the sharpest possible form, i.e., the mean magnetization in the x-y plane remains “frozen” at its initial value, contradicting the “conventional wisdom” which views mean-field theory as the “long-range limit” of interactions of infinite range. Thus, the phenomena of energetic and dynamic stability violate, in general, the universality which was proved, as discussed in the previous subsection, for the critical exponents. In particular, nonequilibrium mean-field theory fails to describe these phenomena, by Theorem 3.5.
Remark 3.2.
The rate of decay of the states $\omega_{p,q}$ also permits to distinguish between long-range and short-range interactions: for various short-range interactions (of exponential decrease), exact results exist which demonstrate that the decay is polynomial ([Em66], [Ra70]).
4 Conclusion and open problems
In [Gli76], J. Glimm remarks: “The importance of universality in mathematics is that indicates the existence of a general theory, whose qualitative (and quantitative) features describe a broad range of phenomena”.
In this paper, however, we analysed some aspects of just one universality class, the Ising universality class defined by the equations (1.2.1), (1.2.4), and (1.2.5), and most results concerned only the special Ising case $\delta=0$. This class also includes Euclidean quantum field theory [GJ81].
The main objects in a future general theory might be microscopic and macroscopic states of infinite systems (this point relates to the second notion of phase transition discussed in Subsection 1.1), and their mutual relationship. The “discontinuities”, mentioned by Glimm as the “essential difference in the analysis over finite vs. infinite dimensional spaces” ([Gli76], p.674), which manifest themselves as singularities in the free energy in the thermodynamic limit, might appear also, in this extended view, as changes in basic structural properties of the states of infinite systems, for instance in the change from pure to mixed in the theory of irreversibility in ([Wre20], [Wre23]), or in the passage from mixed to pure states in the theory of superselection sectors originating from interactions with the environment ([NT99], [NW14], see also [Wre23]). In the present paper, the “discontinuity” also arose from taking the limit (1.4.7) in Definition 1.4. Indeed, starting from a (microscopic) equilibrium (KMS) state $\omega_{T}$, the thereby defined (macroscopic) fluctuation state $\omega_{T}^{F}$ may also exhibit basic structural changes with regard to $\omega_{T}$.
One of the main aspects of the mutual relationship between $\omega_{T}$ and $\omega_{T}^{F}$ concerns the relation between the energy gap, for $T=0$, or the rate of clustering of the ergodic components of $\omega_{T}$, for $0\leq T\leq T_{c}$, and certain properties of $\omega_{T}^{F}$; if there is an energy gap at $T=0$, or the clustering is summable ($l^{1}$), at $0<T<T_{c}$, the fluctuations characterizing $\omega_{T}^{F}$ are normal; if the clustering is not summable ($T=T_{c}$), $\omega_{T}^{F}$ has anomalous fluctuations. This was shown in Theorem 2.5 for $\nu=2$ and for $\nu=3$ under the Standard Assumption (2.3.24). The above connections refer only to the connected $2$-point Wightman functions.
In the case of s.s.b. of a continuous symmetry (e.g., $\left|\delta\right|=1$ in (1.2.1)), the absence of a gap has been proved in the case of finite range interactions ( [WFW87], [LPW81]) and absence of $l^{1}$- clustering in [LPW81]. Thus, Criterion A does not apply. For long-range interactions, such as the BCS model, s.s.b. of a continuous symmetry is compatible with an energy gap (see [We71] and references given there).
The basic element of our framework, following previous work by one of us (M.R.) in [MR02], was to emply a smooth version of fluctuation operators, better adapted to Fourier transform methods. Indeed, a distinction between the critical state and the other ergodic equilibrium states is naturally formulated for the connected two-point Wightman function, in terms of its Fourier transform (2.3.14) which is stronger than the condition of non-$l^{1}$- clustering, but allows a considerable progress towards a proof of the Standard Assumption (Theorem 2.4). Its analogue for the higher ($r\geq 4$)-connected Wightman functions – the Scaling Assumption C ((2.4.2) with $\rho^{{}^{\prime}}>0$ in (2.4.3)) – yields a proof of the quasi-free nature of the fluctuation state $\omega_{T_{c}}^{F}$. Since the corresponding equilibrium state $\omega_{T_{c}}$ is an interacting state, this is a third of the basic structural changes discussed in the previous paragraph. A different structural change (see Remark 2.6) was shown to occur in the passage from finite to infinite volume due to Proposition 1.5, which demonstrates that for any finite volume the fluctuation state is, for any temperature, an interacting state. This proposition is a direct corollary of the important work of De Coninck and Gottal [dCG84], which relies on the property of infinite divisibility of the reciprocal of the characteristic function (for finite volume). A further consequence of their results, Proposition 1.6, was used in the proof of Theorem 2.6.
Several aspects of dynamics were shown to be relevant to this general framework. Firstly, a physically intuitive explanation of the quasi-free property of the fluctuation states for all temperatures (given in Remark 1.7 was suggested by the work of Wehrl [We71] on the spin waves in the BCS model, who showed that the fluctuation states do not satisfy the (dynamical, equilibrium) KMS condition. Since the BCS model is a prototype of long-range interactions, and the quasi-free property of the fluctuation state was also proved for short-range interactions in Theorem 2.6, this “non-equilibrium behavior” of the fluctuations does not, apparently, distinguish between the two main universality classes (long-range and short-range). It is therefore natural to investigate what happens with the other macroscopic observables – the intensive observables – in non-equilibrium situations, and whether there a distinction between the two classes may arise. We were able to investigate this question in Subsection 3.2. The two universality classes -long-range and short-range – are known to exhibit different rates of decay of unstable states (see Remark 3.2). But, in addition, a more significant aspect was shown for the first time: in the long-range case, different rates of fall-off of the potential yield different rates of decay, i.e., “dynamical critical indices” violate universality. These results, proved in the Dyson model in Subsection 3.2 are corollaries of the recent theorem of Albert and Kiessling [AlKi17] on the Cloitre function.
The Bleher-Sinai theory, via hierarchical model, is relevant to the description of the universality classes in the Ising model. It correctly describes the Ising model with long-range interaction and the upper critical dimension of the Ising model [ADC21] and is, in fact, the only way to understand the remarkable results on universality reviewed in Subsection 3.2. Their nonlinear equation for the block spin distribution does have a (non-unique, not normal) Gaussian solution, shown to be correct for the Ising model with short-range interactions by Theorem 2.6. It thereby correctly accounts for the anomalous fluctuations, the major result of physical significance.
The following remarks concern the breakdown of universality in certain models which are not in the Ising universality class, but are nevertheless very illuminating regarding the great complexity of the universality issue.
The result of [LPW81] was generalized by Martin [Mar82] in an optimal way to include several types of decay of the interaction potential, in particular the case of dimension $\nu=2$, providing an extension of the Hohenberg-Mermin-Wagner theorem ([H67], [MeW66]). The latter theorem precludes a first-order phase transition (i.e., with nonzero spontaneous magnetization) in n-vector models, with $SO(n)$ symmetry. van Enter and Shlosman [vES02] showed, however, that, in a model with sufficiently nonlinear interaction potentials which satisfies the Hohenberg-Mermin-Wagner theorem, and therefore no spontaneous magnetization takes place, there is a first-order transition between two non-magnetized states. In three or more dimensions the result is that the ordinary XY model has a critical point separating the non-magnetized high temperature regime from the magnetized low temperature regime: in the non-linear case this transition is first-order.
Another non-universal example is the 3-state Potts model. in dimension two it has an ordinary critical point. If the range is large but finite, the transition becomes first-order. This means that the same potts permutation symmetry is broken, both models are finite-range but the behavior is again non-universal ([GM07]).
Finally, it may be therefore hoped that extensions of the Bleher-Sinai theory to hierarchical models with continuous symmetries (see [MCG08] and references given there) may bring further insights on the basic open problem of what should replace Criterion A in the case of models with s.s.b. of continuous symmetries. In spite of the fact that most magnetic materials are anisotropic, the universality class of the isotropic model (e.g., $\delta=\pm 1$ in (1.2.1)) is certainly not empty![Pel13]. This is one of the major open problems, whose satisfatory solution would be a further hint that the general theory conjectured in this section really exists.
Appendix A “Ising universality class”
The result in this appendix provides additional support to the expression i the title, which is used in connection with Theorem 1.2, showing that Ising domination occurs in connection with the property of existence of an energy gap. It is contained in [vHBW84], but this reference contains several (sign) misprints which may obscure understanding, and omits certain details. We therefore include a complete proof.
Let
$H_{\Lambda}=-1/2\sum_{x,y\in\Lambda}\left(J_{3}(x,y)(\sigma_{x}^{3}\sigma_{y}^{3}-S^{2})+H_{XY}(x,y)\right)$ (A.0.1)
where
$H_{XY}(x,y)\equiv J(x,y)\left(\sigma_{x}^{1}\sigma_{y}^{1}+\sigma_{x}^{2}\sigma_{y}^{2}\right)$ (A.0.2)
and
$|J_{3}(x,y)|=-J_{3}(x,y)\geq|J(x,y)|$ (A.0.3)
Above, and in contrast with the main text, $\sigma^{i},i=1,2,3$ will denote the three component matrices corresponding to a general spin $S$.
The main, elementary, inequality was called in [vHBW84] “classical domination”, but we prefer to call it “Ising domination”, in connection with Remark 1.3:
Theorem A.1 (Ising domination).
$S^{2}-\sigma_{x}^{3}\sigma_{y}^{3}\geq\pm\left(\sigma_{x}^{1}\sigma_{y}^{1}+\sigma_{x}^{2}\sigma_{y}^{2}\right)$ (A.0.4)
i.e., the $3$-terms in (A.0.1) dominate the $XY$ terms, whether ferro- or antiferromagnetic.
Proof. We first take the plus sign in (A.0.4), and prove that
$S^{2}\geq\vec{\sigma}_{x}\cdot\vec{\sigma}_{y}$ (A.0.5)
To prove (A.0.5), recall that $\vec{\sigma}\equiv\vec{\sigma}_{x}+\vec{\sigma}_{y}$ is a vector sum of two quantum spins of magnitude $S$, and has, therefore, magnitude $L\in[0,2S]$. Equation (A.0.5) is thus equivalent to the equation $\left|\vec{\sigma}_{x}+\vec{\sigma}_{y}\right|^{2}\leq 2S(2S+1)$, but this is evident because $0\leq L\leq 2S$.
In order to prove (A.0.4) with the minus sign, divide $\Lambda$ into two sublattices, defined by the nearest neighbors of a point, and perform the unitary transformation $\sigma_{x}^{1}\rightarrow-\sigma_{x}^{1}$ and $\sigma_{x}^{2}\rightarrow-\sigma_{x}^{2}$, together with $\sigma_{x}^{3}\rightarrow\sigma_{x}^{3}$, for $y$ in one sublattice, i.e., a rotation of $\pi$ around the $3$- axis in one sublattice. This concludes the proof of (A.0.4), and thereby of Theorem A.1.
$\Box$
Corollary A.2.
Ground state and energy gap for model (A.0.1), with general spin $S$, long-range interactions, both signs of in the $XY$ term corresponding to ferro or antiferromagnetic couplings, and without requiring translation invariance.
The tensor products of the states $|+S)_{x}$ and $|-S)_{x}$, for all $x\in\Lambda$, are ground states of $H_{\Lambda}$, given by (A.0.1), (A.0.2) and (A.0.3) with eigenvalue zero. If
$|J_{3}(x,y_{0})|-|J(x,y_{0})|\geq\delta>0\text{ for all }x\in\Lambda\text{ andany }y_{0}\text{ fixed }\in\Lambda$ (A.0.6)
then, there is an energy gap in the spectrum of the physical Hamiltonians, i.e, any one of the (positive) Hamiltonians associated to one of the given product states, in the thermodynamic limit.
Proof. By (A.0.1), (A.0.2), (A.0.3), and (A.0.4),
$H_{\Lambda}\geq 1/2\sum_{x,y\in\Lambda}\left(|J_{3}(x,y)|-|J(x,y)|\right)\left(S^{2}-\sigma_{x}^{3}\sigma_{y}^{3}\right)$ (A.0.7)
The fact that the ground states are the two product states $\left|+S\right)_{x}$ and $\left|-S\right)_{x}$, for all $x\in\Lambda$, follows from (A.0.7).
Consider, now, the Hamiltonian $H_{\Lambda}^{+}$, defined by imposing $|+S)_{x}$ boundary conditions at each point $x$ of the boundary of $\Lambda$. By (A.0.6) and (A.0.7), it follows that in the orthogonal complement of the tensor product of the states $|+S)_{x}$, if just one spin is flipped (at some fixed $y_{0}$, now including the boundary) the following inequality holds:
$H_{\Lambda}^{+}\geq\delta S>0$ (A.0.8)
The same result follows for the Hamiltonian $H_{\Lambda}^{-}$, defined by imposing $|-S_{x})$ boundary conditions:
$H_{\Lambda}^{-}\geq\delta S>0$ (A.0.9)
Since (A.0.8), (A.0.9) are uniform in $\Lambda$, the rest of the assertions of Corollary A.2 follows. It is to be noted that, as remarked in [KN97], the energy gap in the thermodynamic limit is a well-defined notion only with respect to one of the superselection sectors, because the corresponding states are disjoint.
$\Box$
Remark A.1.
Corollary A.2 was, for $\nu=1$, and nearest-neighbor interactions, considerably extended by Koma, Nachtergaele, Spitzer and Starr ([KN97], [KN98], [KNS01], [NSS07]) to include kink boundary conditions and thereby proving the existence of domains. Surprisingly, the value of the energy gap is not affected by the existence of a domain wall, as first proved in [KN97] as a consequence of a quantum-group $SU_{q}(2)$ symmetry of the model.
Appendix B Possible non Gaussian behavior at $\nu=4$
A simple general probabilistic approach has been proposed by Newman [Ne79], of which we present a short review.
It is complementary to the present paper because it concerns the dimension $\nu=4$ in the critical case, or $\nu=3$ in the tricritical case.
Let
$\tau_{n}^{2}\overset{\mathrm{def.}}{=}\mathrm{var}(\sigma_{\Lambda_{n}})=\mathbb{E}\sigma_{\Lambda_{n}}^{2}-\left(\mathbb{E}\sigma_{\Lambda_{n}}\right)^{2}$ (B.0.1)
be the variance of the block spin variable with respect to the equilibrium measure (1.1.1) with non-necessarily symmetric a priori measure $\mu_{0}(\sigma)$ on $\mathbb{R}$ and let
$f_{n}(z)=\frac{1}{\left|\Lambda_{n}\right|}\log\mathbb{E}\exp\left(z\sigma_{\Lambda_{n}}\right)~{}.$ (B.0.2)
Tipically $\sigma_{\Lambda_{n}}=\sum_{x\in\Lambda_{n}}\sigma_{x}$ represents some extensive quantity such magnetization, energy or number of particles contained in a finite box $\Lambda_{n}\subset\mathbb{Z}^{v}$ of volume $\left|\Lambda_{n}\right|=n^{\nu}$ and the expectation $\mathbb{E=E}_{\beta,J,\Lambda_{n}}$ with respect to (1.1.1) omits for simplicity the dependences on its parameters and domains. Associated with $\sigma_{\Lambda_{n}}$, $z$ is the intensive thermodynamic variable such as external magnetic field, inverse temperature or the negative of the chemical potential. Free or periodic conditions are imposed in the Hamiltonian of (1.1.1) at the boundary $\delta\Lambda_{n}$ (with magnetic field omitted and possibly long range interactions $J_{xy}$, $x,y\in\Lambda_{n}$).
The following are hypotheses to be assumed accordingly:
  1. A.
    $\exists\varepsilon>0$ so that $f_{n}(z)\longrightarrow f(z)$ exists for $z\in\left(-\varepsilon,\varepsilon\right)$ and $f(z)$ is a real convex function on $\left(-\varepsilon,\varepsilon\right)$ with $f(0)=0$;
  2. B.
    $f^{\prime}(0)$ exists;
  3. C.
    $\left.\mathbb{E}\sigma_{\Lambda_{n}}\right/\left|\Lambda_{n}\right|$ is independent of $n$;
  4. D.
    $\left|\Lambda_{n}\right|\longrightarrow\infty$
  5. E.
    $\sqrt{\mathrm{var}(\sigma_{\Lambda_{n}})}>0$
  6. F.
    $\sqrt{\mathrm{var}(\sigma_{\Lambda_{n}})}\rightarrow\infty$
  7. H.
    $f_{e}(z)\equiv f(z)+f(-z)$ satisfies
    $\sup_{z>0}\lim\sup\limits_{t\rightarrow 0}\frac{f_{e}(tz)}{z^{p}f_{e}(t)}<\infty$
    for some real number $p$;
  8. I.
    $f_{n}(z)\leq f(z)$, $\forall n$ and $z\in\left(-\varepsilon,\varepsilon\right)$.
Weak convergence (or convergence in distribution) of a random variables sequence $X_{n}$, $n\geq 1$, means:
$X_{n}\overset{\mathrm{w}}{\longrightarrow}X\Longleftrightarrow\mathbb{E}g(X_{n})\longrightarrow\mathbb{E}g(X)$
for all bounded continuous functions $g$. Equivalently: $\mathbb{P}(X_{n}\leq\xi)\longrightarrow$ $\mathbb{P}(X\leq\xi)$, $\forall\xi\in\mathbb{R}$ so that $\mathbb{P}(X=\xi)=0$. Write (for $\tau_{n}>0$)
$X_{n}=\frac{\sigma_{\Lambda_{n}}-\mathbb{E}\sigma_{\Lambda_{n}}}{\tau_{n}}\ .$ (B.0.3)
Since $\mathbb{E}X_{n}^{2}=1$, it follows that $\left(X_{n}\right)_{n\geq 1}$ is weakly sequentially compact hence, every subsequence of $\left(X_{n}\right)_{n\geq 1}$ has a weakly convergent subsequence.
When B. is satisfied, $f_{\pm}(z)=f(\pm z)\mp f^{\prime}(0)z$ are non-negative, nondecreasing function on $\left(-\varepsilon,\varepsilon\right)$ with $f_{\pm}(0)=0$ and the following equality $f_{e}(z)=f_{+}(z)+f_{-}(z)$ holds.
Theorem B.1 (main).
  1. 1.
    Suppose A. B. and D. hold. Then
    $\frac{\sigma_{\Lambda_{n}}}{\left|\Lambda_{n}\right|}\overset{\mathrm{w}}{\longrightarrow}f^{\prime}(0)\ .$
  2. 2.
    Suppose A., B., C., E. and I. hold. Then
    $\lim\inf_{n\rightarrow\infty}~{}\left|\Lambda_{n}\right|f_{e}\left(z/\tau_{n}\right)>0\qquad\text{for any \quad}z\neq 0\ .$
  3. 3.
    Suppose A., B., C., E., I. and
    $\lim\sup_{n\rightarrow\infty}~{}\left|\Lambda_{n}\right|f_{e}\left(z/\tau_{n}\right)<\infty\qquad\text{for some \quad}z\neq 0$ (B.0.4)
    hold. Then
    $\lim\inf_{n\rightarrow\infty}~{}\left|\Lambda_{n}\right|f_{\pm}\left(z/\tau_{n}\right)>0\qquad\text{for any \quad}z\neq 0\ .$
    Moreover, if $\lim\sup_{n\rightarrow\infty}$ $\tau_{n+1}/\tau_{n}<\infty$ is assumed in addition, then
    $0<\lim\inf_{z\rightarrow 0}\frac{f_{+}(z)}{f_{-}(z)}<\lim\sup_{z\rightarrow 0}\frac{f_{+}(z)}{f_{-}(z)}<\infty~{}.$
  4. 4.
    Suppose A., B., C., F., H. and I. hold and assume that (B.0.4) is also valid. Then $p\in\left[1,2\right]$ and any limit $X$ of the variable $X_{n}$ given by (B.0.3) satisfies (with $c>0$)
    $\mathbb{E}X=0\qquad\text{and}\qquad\mathbb{E}X^{2}=1\ ;$ (B.0.5)
    $\mathbb{E}\exp(tX)\leq\exp\left(\left|ct\right|^{p}/p\right)\ ,\qquad t\in\mathbb{R}\ ;$ (B.0.6)
    $\mathbb{P}\left(X<-\xi\right),~{}\mathbb{P}\left(X>\xi\right)\leq\exp\left(-\left(\xi/c\right)^{q}/q\right)\ ,\qquad\xi\geq 0\ ,$ (B.0.7)
    where $q\in[2,\infty)$ satisfies $11/p+1/q=$ 1. Thus, if $p\neq 2$, $X$ is non-Gaussian.
See Sec. III of [Ne79] for the proof of assertions 1-4 of Theorem B.1 and Sec. II of [Ne79] for certain classes of models for which the assertions 1-4 are valid.
Remark B.1.
Hypotheses A., B., C., F. and I., which are assumed to hold in order to imply assertions 1-4, are in fact valid at any respectable critical point. Hypothesis I. can actually be proven by FKG (or GKS) inequality, together with translation invariance and reflection positivity, provided A. is assumed valid – the ingenious proof of Proposition 2.5 in [Ne79] is recommended to the reader.
Observe that
$\tau_{n}^{2}=\mathrm{var}(\sigma_{\Lambda_{n}})=\sum_{x,y\in\Lambda_{n}}\left(\mathbb{E}\sigma_{x}\sigma_{y}-\mathbb{E}\sigma_{x}\mathbb{E}\sigma_{y}\right)\ .$
If
$\mathbb{E}\sigma_{x}\sigma_{y}-\mathbb{E}\sigma_{x}\mathbb{E}\sigma_{y}=F_{n}(x-y),$ (B.0.8)
holds, whether by periodic conditions or FKG, then define
$G(R)=\sum_{\left|x\right|\leq R}F_{n}(x)~{}.$
Proposition B.2.
Suppose (B.0.8) holds with $F(x)\geq 0$ for all $x\in\Lambda_{n}\subset\mathbb{Z}^{\nu}$ and $F(0)=0$. Then, there exist $c_{1},c_{2}>0$ and $K_{1},K_{2}<\infty$ such that (with $\left|\Lambda_{n}\right|=n^{\nu}$)
$K_{1}n^{\nu}G(c_{1}n)\leq\tau_{n}^{2}\leq K_{2}n^{\nu}G(c_{2}n)$ (B.0.9)
holds for all $n$.
The following Corollaries of assertions 1-4 of Theorem B.1 establish a hyper-scaling inequality involving critical exponents and provide sufficient conditions for the scaling limit to be non-Gaussian.
Corollary B.3 (Buckingham-Gunton inequality).
Let $p$ and $\eta$ be defined by
$p$ $=$ $\lim\inf_{z\leftarrow 0+}\frac{\log f_{e}(z)}{\log z}$
$\frac{\nu+2-\eta}{2}$ $=$ $\lim\sup_{n\rightarrow\infty}\frac{\log\tau_{n}}{\log n}~{}.$ (B.0.10)
Then $p\geq 1$ and
$2-\eta\leq\frac{2-p}{p}\nu~{}.$ (B.0.11)
Proof. Since $f_{e}^{\prime}(0)=0$ and $\lim_{z\searrow 0}\log f_{e}(z)/\log z=zf_{e}^{\prime}(z)/f_{e}(z)=1$ by l’Hopital and hypothesis B., there is a subsequence $z_{m}\rightarrow 0+$ satisfying $\log f_{e}(z_{m})/\log z_{m}>1-\varepsilon$ for some $\varepsilon>0$ or, equivalently, $f_{e}(z_{m})>z_{m}^{1-\varepsilon}$ (see the first equation (B.0.10)) that contradicts $f_{e}(z_{m})/z_{m}\rightarrow 0$. So, we have $p\geq 1$. It follows from (B.0.10) and assertion 2. of Theorem B.1 that
$0<\left|\Lambda_{n}\right|f_{e}\left(z/\tau_{n}\right)<n^{\nu}\left|\frac{z}{\tau_{n}}\right|^{p^{\prime}}<n^{\nu}\left|\frac{z}{n^{r^{\prime}}}\right|^{p^{\prime}}$
holds for any $p^{\prime}<p$ and $r^{\prime}<\left(\nu+2-\eta\right)/2$. Hence
$\nu-p\frac{\nu+2-\eta}{2}\geq 0\ .$ (B.0.12)
which is equivalent to (B.0.11).
$\Box$
Remark B.2.
The critical exponent $p$ for the Ising model is $p=1+1/\delta$ where $\delta$ is the exponent for the magnetizations as a function of external field at criticality: $f^{\prime}(h)\approx\left|h\right|^{1/\delta}$ at $\beta=\beta_{c}$, as $h\searrow 0\!$. Equation (B.0.11) thus reads
$2-\eta\leq\frac{\delta-1}{\delta+1}\nu$ (B.0.13)
The classical exponents for the Curie-Weiss model are $\delta=3$ and $\eta=0$. Exponent $\eta=0$ holds for hierarchical models, including its $N$ vector version. Replacing the classical exponents into (B.0.13) the inequality turns into an equality at dimension $\nu=4$. For short-range Ising models it is believed, and recently some of the statements have been proved in [ADC21], that a critical dimension $\nu_{c}$ ($\nu_{c}=4$) exists such that for $\nu\geq\nu_{c}$ the scaling limit of the block spin variable is Gaussian, with classical value exponents, and for $\nu<\nu_{c}$ the scaling limit is non-Gaussian and $\delta$ and $\eta$ may not have their classical values. There are some examples of non-Gaussian scaling limit for hierarchical model whose explicit construction, for instance [Fe87, Zu94, BAHM02, MCG08], holds – or allowed to be extended – in dimensions $42<\nu<$ 4. It would be interesting to check whether the hyper-scaling predictions or other consequences in the following Corollaries could be verified for such models.
Corollary B.4.
Suppose that $p$ and $\eta$ defined in Corollary B.3 satisfy (B.0.13) as an equality
$2-\eta=\frac{\delta-1}{\delta+1}\nu\ ,$ (B.0.14)
and there are constants $\theta$ and $\psi$ such that we have
$f_{e}(z)$$\leq$ $K\left|z\right|^{p}\left|\log\left|z\right|\right|^{\theta}\ ,\qquad\text{for }z\ \text{small}$
$\tau_{n}^{2}$$\geq$ $\varepsilon n^{\nu+2-\eta}\left|\log n\right|^{\psi}\ ,\qquad\text{for }n\ \text{large}$ (B.0.15)
for some $\varepsilon$, $K\in\left(0,\infty\right)$. Then $\psi\leq 2\theta/p$.
Suppose that $\psi=2\theta/p$ and, in addition (with (B.0.12) as an equality the lower bound follows from (B.0.15)),
$\varepsilon n^{2\nu/p}\left|\log n\right|^{2\theta/p}\leq\tau_{n}^{2}\leq K^{\prime}n^{2\nu/p}\left|\log n\right|^{2\theta/p}$ (B.0.16)
for some $K^{\prime}<\infty$. Then $p=1+1/\delta\in\left[1,2\right]$ and any limit $X$ of $X_{n}=(\sigma_{\Lambda_{n}}-\mathbb{E}_{\Lambda_{n}}\sigma_{\Lambda_{n}})/\tau_{n}$ in distribution satisfies (B.0.5), (B.0.6) and (B.0.7). Hence, $X$ is non-Gaussian if $p\neq 2$.
The next corollary is intended to deal with ordinary critical point in $\nu=4$ dimensions and tricritical point in $\nu=6$ dimensions. It immediately follows from Corollaries B.3 and B.4 that
Corollary B.5.
Assume that (B.0.9) and
[…]
and
[…]
hold for some $\varepsilon$, $K$, $K^{\prime}\in\left(0,\infty\right)$ and $\theta$, $\theta^{\prime}$, $\psi$, $\psi^{\prime}\in\left(-\infty,\infty\right)$. Then, if $\psi>3\theta/2$ and $\psi^{\prime}>5\theta^{\prime}/3$ any limit $X$ of $X_{n}=(\sigma_{\Lambda_{n}}-\mathbb{E}_{\Lambda_{n}}\sigma_{\Lambda_{n}})/\tau_{n}$ in distribution satisfies (B.0.5), (B.0.6) and (B.0.7) for some $c>0$ with $p=4/3$ and $q=4$ in the former and with $p=6/3$ and $q=6$ in the latter cases.
Remark B.3.
According to Lee-Yang theory as it is modernly stated in [Ne75], a random variable $X$ is said of Newman’s type $\mathcal{L}$, if i. there exist some constants $c,C>0$ such that $\left|\mathbb{E}\exp(zX)\right|\leq C\exp\left(c\left|z\right|^{2}\right)$ holds for all $z\in\mathbb{C}$; and ii. the moment generating function $\mathbb{E}\exp(zX)$ is an even function with only pure imaginary zeros. When $X=\sum_{i\in\Lambda_{n}}\sigma_{i}=\sigma_{\Lambda_{n}}$, the so called Lee-Yang property, satisfied for Dyson hierarchical Ising model [HHW01], see also [CE78], hierarchical $O(N)$ sigma model [Ko88, Zu94, Wa04] and hierarchical spherical model [BAHM02, MCG08], in combination with the renormalization group (RG) machinery, has allowed to deal with both the strong and weak coupling regimes [HHW01, Wa04, MCG08], starting the RG dynamics from their respective original “a priori” measures, evolving towards the fixed point at the critical point. Although the original goal of Lee-Yang program was to understand phase transitions directly from the distribution of Lee-Yang zeroes, rigorous results are very few ([CJN23, CJN21] and references therein) and restrict to specific models ([CJN23, MCG08] and references therein).
Acknowledgements
We would like to thank Bruno Nachtergaele for correspondence on the subject of Appendix A and Aernout van Enter for useful remarks.
We are, above all, grateful to the referee, who advised us to eliminate a large part of irrelevant material, and concentrate only in the Ising universality class. This led us to a streamlined version, in which the open problems are clearly separated from the true subject of the paper. We also thank him for several important remarks, which contributed to render the discussion completely precise, and considerably enriched the paper with observations, mostly regarding the concept of universality.